### v1

# Improving Sequential Recommendations with LLMs

Abstract.
The sequential recommendation problem has attracted considerable research attention in the past few years, leading to the rise of numerous recommendation models. In this work, we explore how Large Language Models (LLMs), which are nowadays introducing disruptive effects in many AI-based applications, can be used to build or improve sequential recommendation approaches. Specifically, we design three orthogonal approaches and hybrids of those to leverage the power of LLMs in different ways. In addition, we investigate the potential of each approach by focusing on its comprising technical aspects and determining an array of alternative choices for each one. We conduct extensive experiments on three datasets and explore a large variety of configurations, including different language models and baseline recommendation models, to obtain a comprehensive picture of the performance of each approach.
Among other observations, we highlight that initializing state-of-the-art sequential recommendation models such as BERT4Rec or SASRec with embeddings obtained from an LLM can lead to substantial performance gains in terms of accuracy. Furthermore, we find that fine-tuning an LLM for recommendation tasks enables it to learn not only the tasks, but also concepts of a domain to some extent. We also show that fine-tuning OpenAI GPT leads to considerably better performance than fine-tuning Google PaLM 2. Overall, our extensive experiments indicate a huge potential value of leveraging LLMs in future recommendation approaches. We publicly share the code and data of our experiments to ensure reproducibility.  [^1] https://github.com/dh-r/LLM-Sequential-Recommendation
Recommender Systems, Large Language Models, Sequential Recommendation, Evaluation
ccs: Information systems Recommender systems
1. Introduction
The sequential recommendation problem considers a sequence of past user interactions in order to predict the next user interest or action. The problem’s wide application in many popular domains, such as next-purchase prediction (Jannach et al., 2015a), next-track music recommendation (Chen et al., 2012), and next Point-of-Interest suggestions for tourism (Lim et al., 2015), has sparked substantial research interest in the recent years. This had led to the inception of novel algorithmic approaches (Hidasi et al., 2016; Ludewig and Jannach, 2018; Kang and McAuley, 2018; Sun et al., 2019), including approaches that utilize side information about the items, such as an item’s category (Xu et al., 2022; Lai et al., 2022).
Lately, the advent of Large Language Models (LLM), which in an auto-regressive mode repeatedly complete a sequence until they fulfil a prompt task (Radford et al., 2018), has opened a fruitful research direction for sequential recommendation. In fact, LLMs based on Generative Pretrained Transformers (Radford et al., 2019) are only the latest of a stream of innovations in Natural Language Processing (NLP) that inspired the birth of new sequential recommendation models. Notable examples are GRU4Rec (Hidasi et al., 2016), SASRec (Kang and McAuley, 2018), and BERT4Rec (Sun et al., 2019), which were influenced by the Gated Recurrent Unit (GRU) (Cho et al., 2014), the transformer architecture (Vaswani et al., 2017), and BERT (Devlin et al., 2019), respectively. The key contribution of LLMs to the sequential recommendation problem is a semantic model of real world concepts and their representation in natural language sequences. This consideration is neglected by typical sequential recommendation models (Kang and McAuley, 2018; Sun et al., 2019).
The rapidly growing interest for involving LLMs in recommender systems clusters around cross-domain (Zhang et al., 2021a; Hou et al., 2023b), domain-specific (Bao et al., 2023; Zhang and Wang, 2023), and multi-modal (Wu et al., 2022; Wang et al., 2022) recommendations. Within the scope of recommendation approaches, LLMs are involved either by providing item embeddings (Hou et al., 2023a) or by responding to a recommendation task conveyed in a prompt, which may contain zero (Ding et al., 2022) or a few examples (Geng et al., 2022) to enrich the context of the task available to an LLM. Alternatively, an LLM can be fine-tuned to learn how to perform better for a specific recommendation task (Hou et al., 2022). Regarding the nature of recommendation tasks posed to LLMs, they are typically used to generate (Li et al., 2023f) or rerank (Hou et al., 2023b) recommendations. Altogether, the current state of research is summarized in the following surveys (Lin et al., 2023; Wu et al., 2023). Given the currently early stage of this research avenue, the prospect for the various recommendation approaches involving LLMs remains largely unexplored.
In this work, [^2] The present paper substantially extends our previous work on leveraging LLMs for the sequential recommendation problem (Harte et al., 2023). We detail our introduced extensions in Section 2. we design and research the potential of three orthogonal approaches and hybrids of those that leverage LLMs for sequential recommendations. We decompose each approach into its constituent technical aspects, identify or devise alternatives for each aspect, and investigate the impact of choices to the approach’s performance. Finally, we thoroughly tune and evaluate all approaches under the same testbed across three datasets of different domains and characteristics. These datasets are Amazon Beauty from the e-commerce beauty domain, a proprietary Delivery Hero dataset from the e-commerce darkstore domain, and Steam from the gaming domain.
In the first approach (LLMSeqSim), we retrieve a semantically-rich embedding from an existing LLM for each item in a session, optionally reduce the embedding to a target number of dimensions, and then compute an aggregate session embedding. We use the latter to recommend catalog products with a similar embedding. In this first approach we investigate the following four technical aspects: the choice of LLM to retrieve embeddings from, dimensionality reduction methods, target embedding dimensions, and session embedding computation strategies. In the second approach (LLMSeqPrompt), we fine-tune an LLM with dataset-specific information in the form of prompt-completion pairs and ask the model to produce next item recommendations for a set of test prompts. We explore four different aspects of LLMs, namely models, versions, parameters, and task specifications. Our third approach (LLM2Sequential) consists of enhancing existing sequential models with item embeddings obtained from an LLM. Here, we examine LLM embedding models, dimensionality reduction methods, and sequential models. Finally, we create hybrid recommendation approaches based on the aforementioned ones.
Our work results in the following contributions and insights.
  • We design and research three orthogonal methods and two hybrid approaches of leveraging LLMs for sequential recommendation. In particular, we investigate the technical aspects of each method, determine alternatives for each, thoroughly tune them, and evaluate their impact.
  • Experiments on three datasets, including a proprietary real-world dataset from Delivery Hero, reveal that LLM embeddings boost the performance of different classes of sequential models across all datasets. Specifically, LLM2SASRec and LLM2BERT4Rec, which are implementations of the LLM2Sequential approach, top the leaderboard increasing NDCG@20 on average by 45% on the Amazon Beauty dataset and 9% on the Delivery Hero dataset. Beyond accuracy metrics, LLM2SASRec almost doubles catalog coverage and substantially increases serendipity by 21% over SASRec across all datasets.
  • Both OpenAI GPT and Google PaLM fine-tuned models (LLMSeqPrompt) perform significantly better than their counterpart base models across tasks marking 16% average improvement in terms of NDCG@20. Most importantly, the fine-tuned models exhibit as low as half the number of hallucinations and noticeably higher semantic similarity to actual recommendation options compared to the base models indicating that fine-tuned models can learn not only tasks, but also concepts of a domain to some extent. Finally, GPT 3.5 turbo considerably outperforms PaLM 2 bison by more than 78% on average in terms of NDCG@20 across fine-tuning tasks and datasets.
  • In one of the datasets, Amazon Beauty, the semantic item recommendation model via LLM embeddings (LLMSeqSim) takes the third place and achieves the best MRR score, while a fine-tuned model (LLMSeqPrompt) outperforms GRU4Rec and SASRec. Both results support the potential of LLM-based models for datasets with specific characteristics.
The rest of the paper is organized as follows. Section 2 presents the background and related work. Next, Sections 34,  5, and 6 present the researched recommendation approaches. Then, Section 7 details the experiments and, finally, Section 8 conveys the conclusions of our work.
2. Background & Related Work
The recent developments in LLMs have taken the world by surprise. Models like OpenAI GPT (Brown et al., 2020), Google BERT (Devlin et al., 2019), and Meta LLaMA (Touvron et al., 2023), which employ deep transformer architectures, demonstrate how innovations in NLP can reshape and advance mainstream online activities, such as search, shopping, and customer care. Inevitably, research in recommender systems is significantly impacted by the developments in the area of LLMs as well.
According to recent surveys (Lin et al., 2023; Wu et al., 2023), LLMs are mainly utilized for recommendation problems in two ways: by providing embeddings that can be used to initialize existing recommendation models (Wu et al., 2021; Zhang et al., 2021b; Liu et al., 2021), and by producing recommendations leveraging their inherent knowledge encoding (Kang et al., 2023; Hegselmann et al., 2023; Bao et al., 2023). LLMs as recommendation models can provide recommendations given (i) only a task specification (zero-shot), (ii) one or few examples given inline to the prompt of the task (few-shot), or (iii) after fine-tuning the model’s weights to the task at hand given a set of training examples (Brown et al., 2020). This incremental training process deviates from typical recommendation models, which have to be trained from zero on domain data. In fact, LLMs show early indications of adaptability to different recommendation domains with modest fine-tuning (Hou et al., 2022, 2023a). Finally, LLMs have so far been used for a number of recommendation tasks, such as rating prediction (Li et al., 2023c), item generation (Li et al., 2023f), and reranking (Hou et al., 2023b) across domains (e.g., news (Wu et al., 2021), information retrieval (Liu et al., 2021)).
In this work we explore the potential of using LLMs for sequential recommendation problems (Jannach et al., 2020). In these problems, we consider as input a sequence of user interactions (i.e.a session) $S^{u}=(S^{u}_{1},S^{u}_{2},\ldots,S^{u}_{n})$, where $u$ is a user, $n$ is the length of the sequence and $S^{u}_{i}$ are individual items. The aim is to predict the next interaction of the given sequence. Besides the recent sequential recommendation models mentioned in the introduction (Hidasi et al., 2016; Kang and McAuley, 2018; Sun et al., 2019), in earlier works the sequential recommendation problem has been modelled as a Markov Chain (Garcin et al., 2013) or a Markov Decision Process (Shani et al., 2005). Neighborhood-based approaches, such as SKNN (Jannach and Ludewig, 2017), have also been proposed.
Related pieces of work in the sequential recommendation problem focus on cross-domain (Section 2.1), domain-specific (Section 2.2), and multi-modal (Section 2.3) recommendations.
2.1. Cross-domain LLM-based Sequential Recommendation Models
Early research work regarding LLMs for sequential recommendation problems showed mixed results for zero-shot and fine-tuned recommendations (Zhang et al., 2021a; Geng et al., 2022; Ding et al., 2022; Liu et al., 2023b; Hou et al., 2023b, a; Wang and Lim, 2023; Sileo et al., 2022). The first deep zero-shot generative recommender model named ZESRec (Ding et al., 2022) comprises four unique properties: cold users (i.e., no overlapping users between training and test data); cold items (i.e., no overlapping items between training and test data); domain gap (i.e., training and test data originate from different domains); and no access to target data (i.e., target data are only available at inference time). ZESRec outperforms the zero-shot embedding-KNN and random baselines substantially, and proves beneficial for data-scarce startups and early-stage products. In (Wang and Lim, 2023) and (Zhang et al., 2021a) both GPT (Radford et al., 2018) and BERT (Devlin et al., 2019) perform better than a random baseline in zero-shot sequential recommendations, while GRU4Rec (Hidasi et al., 2016) outperforms the fine-tuned versions of the LLMs. GRU4Rec and other baselines also outperform ZESRec (Ding et al., 2022) for zero-shot recommendations across domains. Conversely, P5 (Geng et al., 2022), which can produce zero-shot or few-shot recommendations without fine-tuning, has been shown to outperform a number of sequential recommendation models and ChatGPT—the least performant of all models (Liu et al., 2023b).
The very recent VQ-Rec model (Hou et al., 2023a) employs a transformer architecture and applies a novel representation scheme to embeddings retrieved from BERT to adapt to new domains. VQ-Rec outperforms a number of sequential recommendation models across datasets of different domains. Chat-Rec (Gao et al., 2023) converts user profiles and past interactions into prompts to build a conversational recommender system based on ChatGPT. The system is effective in learning user preferences and transferring them to different products, potentially allowing cross-domain recommendations. Finally, in both (Hou et al., 2023a) and (Yuan et al., 2023) SASRec with LLM embeddings is shown to improve over SASRec.
To explore the performance limits of text-based collaborative filtering with LLMs, Li et al(Li et al., 2023b) conduct a series of experiments contrasting the former with the dominant identifier-based approach and the trending prompt-based recommendation using ChatGPT. Interestingly, ChatGPT seems inferior than text-based collaborative filtering, while the simple identifier-based approach remains highly competitive in the warm item recommendation setting. However, in the book recommendation tasks (i.e., book rating, user rating, and book summary recommendation), a ChatGPT-like system (BookGPT) developed by Li et al(Li et al., 2023a) achieves promising results, especially in zero-shot or one-shot learning, compared to the classic book recommendation algorithms. Another investigation on the capabilities of ChatGPT in recommender systems (Dai et al., 2023) highlights that the model performs better with list-wise ranking, compared to point-wise and pair-wise. Overall, LLM-based recommenders surpass popularity-based recommenders in the movie, book, and music domains, but not so in the news domain.
2.2. Domain-specific LLM-based Sequential Recommendation Models
In the movie, book, and e-commerce domains, the lightweight TALLRec framework proposed by Bao et al(Bao et al., 2023) aims to adapt LLMs for recommendations by structuring the recommendations as instructions, and tuning the LLMs through an instruction-tuning process. Experiments using the LLaMA model (Touvron et al., 2023) demonstrate that TALLRec significantly advances the LLM capabilities, and is highly efficient even on a single conventional GPU. Similarly, PALR (Yang et al., 2023) employs user/item interactions for candidate retrieval, and a fine-tuned LLM-based ranking model to produce recommendations. Using LLaMA for the LLM, and compared to representative baselines including GRU4Rec (Hidasi et al., 2016) and SASRec (Kang and McAuley, 2018), PALR appears more effective in various sequential recommendation tasks. Furthermore, GPTRec (Petrov and Macdonald, 2023), which is based on the GPT-2 architecture (Radford et al., 2019), employs a novel SVD tokenization algorithm and a Next-K recommendation strategy to generate sequential recommendations, accounting for already recommended items. GPTRec matches the quality of SASRec (Kang and McAuley, 2018) but is more GPU-memory efficient due to its sub-item tokenization. Inspired by search engines, GPT4Rec (Li et al., 2023f) generates a set of search queries based on the item titles from a user’s history, and then recommends items by searching these queries using beam search. Using GPT-2 (Radford et al., 2019) as a language model to generate queries that are used in a search engine using BM25 (Robertson and Zaragoza, 2009) as its score function, GPT4Rec outperforms state-of-the-art models including YouTubeDNN (Covington et al., 2016) and BERT4Rec (Sun et al., 2019). In addition, multi-query generation with beam search increases the diversity of recommended items and the coverage of a user’s interests.
In the news domain, Zhang and Wang (Zhang and Wang, 2023) have recently released a framework for prompt learning (Prompt4NR) to predict whether a user would click a candidate news item. Prompt4NR slightly outperforms BERT-based models, reflecting the usefulness of the prompt learning approach for embedding knowledge in the pre-trained BERT models employed for news recommendation. Similarly, the prompt-based news recommender system (PBNR) by Li et al(Li et al., 2023d) leverages the text-to-text T5 model (Raffel et al., 2020) for personalized recommendations. PBNR is very accurate even with varying lengths of past user interactions, adaptable to new data, and can also satisfy user requirements through human-computer interaction. Finally, Li et al. investigate ChatGPT’s performance in news recommendation (Li et al., 2023e) with respect to personalized recommendations, fairness, and fake news detection.
To enhance content-based recommendation, Liu et al(Liu et al., 2023a) propose a framework named ONCE which combines the deep layers of open-source LLMs (i.e., LLaMA (Touvron et al., 2023)) as content encoders to improve the representation embeddings, and the prompting techniques of closed-source LLMs (i.e., ChatGPT (Liu et al., 2023b)) to enrich the training token data. ONCE shows a substantial improvement of 19.32% compared to the state-of-the-art (NAML (Wu et al., 2019a), NRMS (An et al., 2019), Fastformer (Wu et al., 2019b)), showcasing the synergistic relationship between fine-tuning on the open-source LLMs and prompting on the closed-source LLMs. UniTRec, a unified text-to-text transformer and joint contrastive learning framework for text-based recommendation is proposed by Mao et al(Mao et al., 2023) to better model two-level contexts of user history. The framework reaches state-of-the-art performance on three text-based recommendation tasks, namely news, quotation, and social media post recommendation.
2.3. Multi-modal LLM-based Sequential Recommendation Models
A number of research works combine both textual and visual information to develop multi-modal sequential recommendation models. MM-Rec by Wu et al(Wu et al., 2022) includes a cross-modal candidate-aware attention network that selects relevant historical clicked news and models user interest in candidate news. Experiments suggest that multi-modal news information accelerates news recommendation performance. TransRec (Wang et al., 2022) slightly modifies the popular identifier-based recommendation framework by directly learning from the raw features of the multi-modal items in an end-to-end style, enabling transfer learning without depending on overlapped users or items. TransRec appears to be a generic model that can be transferred to various recommendation tasks. However, it entails a high training cost due to its end-to-end learning method, and can only be applied to image-only, text-only, and image-text scenarios.
To reduce TransRec’s high training cost, Fu et al(Li et al., 2023b) research whether an adapter-based variant of TransRec with two task-specific neural modules inserted into each Transformer block could perform similarly to the original fine-tuned model (Wang et al., 2022). Their findings suggest a comparable performance between the two for both text and image recommendations. To shed light on the modality-based versus identifier-based recommendation architecture dilemma, Yuan et al(Yuan et al., 2023) conduct an empirical investigation tackling various associated questions. They conclude that modality-based recommendation is comparable to identifier-based, and could potentially surpass it given the rapid NLP and computer vision advances. Still, the former is more expensive in training, aligning with TransRec’s findings (Wang et al., 2022).
2.4. Comparison to Previous Paper and Related Work
Our present work builds on top of our previous paper (Harte et al., 2023), where we proposed three orthogonal approaches of leveraging LLMs for the sequential recommendation problem. The first one, LLMSeqSim, obtains from an LLM the embeddings of items in a session, combines them to construct a session embedding, and maps the session embedding to the embedding space of available catalog items in order to produce next-item recommendations. The second approach, LLMSeqPrompt, involves fine-tuning an LLM to generate next-item recommendations. We provided session data in the form of prompt-completion pairs, where the prompt contains all items of each session but the last and the completion contains the last item, that is the ground truth. At prediction time, we asked the fine-tuned model to recommend an item for each session of the test set. Finally, the third approach, LLM2BERT4Rec, entails initializing the embedding layer of BERT4Rec with LLM embeddings to enable BERT4Rec to leverage the embeddings in its training process. Experimental results on the Amazon Beauty and a Delivery Hero dataset showed that this approach improves accuracy in terms of NDCG@20 by 15–20%.
In the meantime, a new body of research work involving LLMs in the sequential recommendation problem has swiftly emerged as we discuss in Sections 2.12.2, and 2.3. With respect to leveraging LLM embeddings for producing recommendations, we identify no similar work to our LLMSeqSim approach. The majority of research efforts target fine-tuning LLMs for various recommendation tasks, and result in mixed conclusions, aligning with our LLMSeqPrompt observations. Researchers have devised a variety of methods including combining closed and open-source LLMs (Liu et al., 2023a), using algorithms from search engines (Li et al., 2023f), utilizing user profiles (Gao et al., 2023), and structuring recommendations as instructions (Bao et al., 2023). Overall, these methods appear mainly suitable for cold-start and data-scarce recommendation problems, and highly depend on the applied domain and dataset (Dai et al., 2023).
Regarding the initialization of neural sequential models with LLM embeddings, the recent approaches presented in (Hou et al., 2023a)(Hua et al., 2023), and (Yuan et al., 2023) differ from our work in particular in terms of the goals they pursue. VQ-Rec (Hou et al., 2023a) targets cross-domain recommendations with a novel item representation scheme, while (Hua et al., 2023) and (Yuan et al., 2023) evaluate whether recommendation models leveraging different modalities perform better than existing recommendation models that rely on item identifiers. In general, there seems to be a common conclusion that enhancing neural sequential models with LLM embeddings advances model performance due to the enriched information contributed by the embeddings (Hou et al., 2023a; Yuan et al., 2023; Zhang and Wang, 2023). As we will notice, this conclusion is also supported by our own results.
In this work we propose multiple and substantial extensions to our previous paper. Overall, the extensions regard (i) a new breed of hybrid approaches born from the above-mentioned approaches and from existing recommendation models, (ii) additional LLM embedding models, (iii) more dimensionality reduction methods, (iv) additional LLMs available for fine-tuning and new versions of those, (v) new fine-tuning task specifications, and (vi) more sequential recommendation models enhanced with LLM embeddings.
These extensions are implemented in three orthogonal plus a new breed of hybrid approaches, shown in Figure 1. Each approach comprises a number of technical aspects for which we devise and investigate alternative choices. We evaluate a large variety of configurations for all approaches on the same experimentation testbed, against the state-of-the-art sequential recommendation models, using three datasets of different domains. In this way, we delve into each of the fundamental approaches regarding the involvement of LLMs in the sequential recommendation problem and assess their potential to advance the existing state of the art and point to promising new avenues for research. Finally, based on encouraging results, we explore hybrid approaches by combining variations of the proposed methods, aiming to use each individual method’s dominant characteristics in an ensemble to improve recommendations further.
Figure 1. Layout of work. Each top-level branch of the tree shows one approach presented in the corresponding section. In brackets we indicate the alternatives we used.
3. LLMSeqSim: Semantic Item Recommendations via LLM Embeddings
Figure 2. Semantic item recommendations via LLM embeddings.
Our goal with this first approach is to explore if recommendations can benefit from a notion of semantic similarity as provided by LLMs. Specifically, we assume that LLM embeddings encapsulate semantic relationships regarding items of many domains based on their massive training data. Therefore, when we encounter a recommendation use case for a particular domain, the LLM embeddings can position items of the domain relative to other items, considering the items’ semantics and their relationships with relevant concepts of the domain. In this way, embeddings can enable recommendations that stem from a deep semantic intuition of the items’ nature. In addition, putting together a sequence of such item embeddings opens a potential for exploring a space of recommendations that complete the sequence in novel ways not envisioned before, as we elaborate in Section 3.4.
In the proposed LLMSeqSim (Sequential Similarity) approach, we leverage LLM embeddings to produce recommendations in six steps. We depict these steps in Figure 2, applied on a food recommendation use case. At the first step we supply metadata of the catalog’s items, such as the names of the products, to an LLM embedding API and retrieve the corresponding embeddings. Then, at the second step, we can optionally reduce the dimensions of the embeddings to improve, if possible, the performance of the recommendation task. Third, we store the embeddings of the item catalog.
When a prompt session arrives, for instance a cart of products, we query the item catalog embeddings using the prompt’s item metadata and retrieve their embeddings at the fourth step. Then, we compute a session embedding for each prompt session in our test set by combining the embeddings of the individual items in the session. As a sixth step, we compare the session embedding to the embeddings of the items in the item catalog using cosine, Euclidean, and dot product similarity. [^3] The choice of the similarity measure did not significantly impact the results. We used cosine similarity. Finally, we recommend the top-$k$ products from the catalog with the highest embedding similarity to the session embedding.
We identified four technical aspects in this processing pipeline, where different implementation alternatives are possible. Below, in Sections 3.13.4, we elaborate our investigation of the different aspects involved in the LLMSeqSim approach.
3.1. Embedding Sources
With regard to LLM sources that provide embeddings in the first step, we explore both OpenAI’s embedding model (text-embedding-ada-002) [^4] https://platform.openai.com/docs/guides/embeddings/second-generation-models as well as Google’s PaLM2-based embedding model (textembedding-gecko@002). [^5] https://cloud.google.com/vertex-ai/docs/generative-ai/embeddings/get-text-embeddings Our motivation for choosing these embedding models is to investigate and uncover commonalities and differences between embeddings provided by diverse and highly performing foundational models.
3.2. Dimensionality Reduction Methods
After retrieving the embeddings, we can directly use them. However, given that the embeddings are of multidimensional nature and especially LLMs utilize a large number of dimensions to model a vast number of concepts in the same space, we take the opportunity to lower the embedding dimensions using a set of dimensionality reduction methods that may improve the performance of the recommendation task.
We applied four different dimensionality reduction methods. The first dimensionality reduction method we used is Principal Component Analysis (PCA), a method that has proved very effective for the initialization of the embedding layer of neural models (Harte et al., 2023). Then, as an alternative to PCA we also tried Linear Discriminant Analysis (LDA), which projects the data to $k-1$ dimensions so that the projected data best fit $k$ classes. We used the most characteristic item metadata to denote classes, such as the name or category of each item depending on the dataset. The third approach was to apply an autoencoder to reduce the dimensions of embeddings, while experimenting also with $\ell 2$ regularization that might further benefit the reduction by driving minor embedding scores to zero. Finally, the fourth approach was a random projection of embeddings (Johnson, 1984) to a lower dimensional space with the premise that the lower space provides a suitable approximation of the original space in terms of distances between the points.
3.3. Number of Reduced Dimensions
In each of the four dimensionality reduction methods we tried different numbers of lower embedding dimensions. The motivation behind this exploration is to discover whether and to which extent dimensionality reduction produces embeddings of higher quality that lead to improvements in recommendation performance.
Our starting dimensions were 768, which is the number of dimensions of Google embeddings, and 1536, which is the number of dimensions of OpenAI embeddings. We tried reducing these to 128 and 512 dimensions to balance the overall hyperparameter search space and the performance impact of the dimensions based on observations from ad-hoc experiments. Prior to the hyperparameter search process, we also explored 256, 64, and less than 64 dimensions manually in ad-hoc experiments, but did not identify any potential for further investigation in low dimensions.
3.4. Session Embedding Computation
[…]
It is possible to use different pooling methods, beyond global average pooling. We therefore try a number of different pooling strategies: (i) the average of the product embeddings, (ii) a weighted average using linear, harmonic, and exponential decay functions depending on the position of the item in the session, and (iii) only the embedding of the last product. [^6] We also tried to create an aggregated session embedding by concatenating the plain product names and then querying the OpenAI embeddings API. This however led to worse results. We provide the exact computation techniques of the weighted average in Table 1, where for each technique we specify the decay computation at each step of a session.
Technique Decay per Step
Constant linear $1/10$
Scaling linear $1/|S|$
Scaling quadratic $1/|S|^{2}$
Scaling cubic $1/|S|^{3}$
Log $1/\log(i+1)$
Harmonic $1/i$
Squared harmonic $1/i^{2}$
Table 1. Computation techniques for weighted averages; $|S|$ stands for the length of session $S$.
4. LLMSeqPrompt: Prompt-based Recommendations by a Fine-Tuned LLM
In this approach, we inject domain knowledge to the collective information that a base LLM incorporates, with the goal of combining fundamental information about the items of the domain encoded in the LLM with domain-specific information conveyed in the fine-tuning process. The composition of these two elements of information exposed in the same highly capable substrate of an LLM paves the path towards potentially advanced recommendations.
To this end, we fine-tune an OpenAI ada [^7] https://platform.openai.com/docs/guides/fine-tuning model, an OpenAI GPT-3.5 turbo [^8] https://platform.openai.com/docs/guides/fine-tuning/what-models-can-be-fine-tuned model and the model text-bison@001999https://cloud.google.com/vertex-ai/docs/generative-ai/models/tune-models, which is based on Google PaLM2. GPT succeeded ada as OpenAI’s recommended base model for fine-tuning, while PaLM2 is a base model created by Google. While there is general consensus on the overall architecture and training process of a base LLM, differences in architectural choices, its configuration, and the training data mean that LLMs can differ significantly. [^10] A discussion of those factors is out of this paper’s scope. Thus, the aforementioned selection of models enables us to compare and contrast two fundamentally different LLMs and also assess the performance evolution and qualitative characteristics of OpenAI models, going from ada to GPT 3.5 turbo.
We fine-tune the models on training samples consisting of a prompt (the input) and a completion (the intended output). First, we take the training sessions and format them as prompt-completion pairs according to the fine-tuning task specification. We present the precise task specifications in Sections 4.14.24.3, and 4.4. We also separate a small subset of the training sessions and provide them to the fine-tuning process as validation set. We supply 500 sessions for validation to OpenAI and 250 sessions to Google PaLM, which is the maximum number of validation cases that the PaLM API allows. Then, we supply the prompt-completion pairs to a model’s fine-tuning API. In all configurations, we fine-tune the model until the validation loss converges in order to achieve the best performance.
After fine-tuning, we provide the prompts of the sessions in the test set to the fine-tuned model and retrieve recommendations. Finally, for generation tasks that may lead to hallucinated recommendations, we map non-existing items back to the item catalog through dot product embedding similarity [^11] We use OpenAI embeddings to resolve hallucinations, which are normalized. Therefore, the dot product is the same as cosine similarity. and obtain the final recommendations.
Besides employing a set of diverse LLMs, we propose four alternative specifications for the fine-tuning task. Our rationale is to examine the behavior and performance of a model on different problem formulations, namely generation of a single recommended item (Section 4.1), generation of a list of item recommendations (Section 4.2), classification (Section 4.3), and ranking (Section 4.4). [^12] Budget and time limitations allowed us to apply the fine-tuning tasks of Sections 4.24.3, and 4.4 to the Beauty dataset only. Generation, classification, and ranking enable an investigation of fine-tuning from less to more restricted settings. Fine-tuning a model on a generation task allows the model to recommend virtually anything. In a classification task, in contrast, a fine-tuned model is presented with a candidate pool of item recommendations presented as categories and selects a top-$k$ subset of those similarly to multi-class classification. Finally, the ranking task further narrows the freedom of the model in that the set of available recommendations is provided in the prompt, allowing the model to focus on the best ranking of those recommendations.
Figure 3. Next-item generation by fine-tuned LLM.
4.1. Fine-tuning Task: Generate Single Recommended Item
In this variant, the prompt is a session, which contains a list of item metadata, e.g., product names, except for the last item, and the completion contains the metadata of the last item in the same session according to the first step in Figure 3. At prediction time, following the fine-tuning process, we repeat a prompt $k$ times ($k=20)$ and collect a single recommendation produced by the model per prompt invocation according to the fine-tune task formulation. We use the tendency of the model to make the same recommendation in different invocations for the same prompt as a proxy of its confidence for the recommendation, collect and deduplicate the recommendations over the $k$ invocations and rank the recommendations by frequency of appearance. The fine-tuned LLM, being a generative model, may also return hallucinated products. To remedy that, in the fifth step, we retrieve the embedding of each deduplicated recommendation using either the OpenAI or the Google embeddings API (this embedding might correspond to a hallucinated product), and finally we take the catalog’s product that is closest in terms of embedding similarity using the dot product.
Repeating the prompt allows us to get a number of different recommendations from the single recommendation returned each time. Alternatively, we can overcome the limitation of a single recommendation per prompt, as we do next.
Figure 4. Recommendation list generation by fine-tuned LLM.
4.2. Fine-tuning Task: Generate List of Item Recommendations
By design, a recommendation model aims to produce a number of desirable options in order to maximize the probability of satisfying a user’s intent at the time. Therefore, we would like to teach an LLM to embrace the role of a typical recommendation model and generate a list of next-item recommendations ranked by order of relevance that is determined by the fine-tuned model.
To convey this task in the fine-tuning process as we depict in Figure 4, we provide in the prompt a session containing a list of items except the last one. In the prompt completion we provide a list of item recommendations that is compiled by a state of the art recommendation model. The rationale is to ask a proven recommendation model to provide recommendations for the training sessions and use those as prompt completions for the fine-tune process. In fact, we can query more than one recommendation model and use the recommendations of the model that achieves the highest NDCG, or another metric, per training session. With enough training sessions, this practice could potentially transfer the knowledge of a powerful ensemble of models to a fine-tuned LLM. [^13] We did not apply this step in this work. Instead, we used LLMSeqSim in the Beauty dataset due to its high performance. Finally, we insert the ground truth item at the top of the list, if not already there, and fine-tune an LLM with the same prompts and the corresponding compiled list of recommendations as completion.
At prediction time, we ask the model to produce recommendations by passing the items of each test session except the last item. The fine-tuned model may return hallucinated recommendations, which we map back to the item catalog following the same method as in Section 4.1.
Figure 5. Multi-class classification of items by fine-tuned LLM.
4.3. Fine-tuning Task: Classify Items for Next Item Recommendations
Asking an LLM to generate responses without explicitly providing the eligible options in the prompt represents the most challenging type of task. The reason is that it forces the model to choose from a virtually infinite set of options often leading to hallucinations. Therefore, we aim to liberate an LLM from the complexity of generating unbounded responses based on its fine-tuning. Instead, we allow it to focus on picking and ranking a list of pertinent item recommendations from an overall pool of items that we pose to the model as classification categories. The intuition behind this formulation is that there exists a modest subset of popular and diverse items of a domain that will please most customers avoiding, at the same time, niche items that receive scant attention.
In this variant, depicted in Figure 5, we state the next item recommendation problem as a multi-class classification task as follows. Prior to the fine-tuning process, we apply a clustering algorithm, namely K-means, to group items to clusters based on the embedding distances of their metadata [^14] We use product names in the Amazon Beauty dataset. As a result, the clusters can be seen as a set of diverse item groups that are potentially representative of the whole item space. We pick a relatively large number of clusters, i.e., 200, to maximize the probability of a good representation of semantically different items by the corresponding clusters. Then we elevate the most popular item of a cluster to represent the group of items in the cluster. In this fashion, we pick an array of popular and diverse candidate items that could represent good-quality recommendations across the total item space. Finally, for each training session, we take the embeddings of the ground truth item and the items that represent the classification categories and select the top-$k$ semantically closest classification categories for each ground truth item.
In the prompts used for fine-tuning, together with each input session of items (except the last item), we include the classification categories and in the corresponding completions we include the top-$k$ classification categories that we computed in the previous step. At prediction time, we ask the model to classify a prompt session to top-$k$ categories, which correspond to next item recommendations. Note that since the candidate recommendations are included in the prompt, there is no need for an extra step to map the recommendations back to the item space. The fine-tuned models respond according to their fine-tuning task specification.
Figure 6. Rank recommendation slate by fine-tuned LLM.
4.4. Fine-tuning Task: Rank Next Item Recommendations
The ranking task enables the fine-tuned model to draw its attention to the ranking of recommendations. By fine-tuning a model to rank recommendations, we aim to investigate the efficacy in the ranking task of an LLM adapted to a particular domain compared to the more open fine-tuning task specifications of Sections 4.14.2, and 4.3.
In this variant, visualized in Figure 6, we fine-tune the models to learn a ranking task where we pass in the prompt a session of items except the last one and a set of recommendations including the ground truth, that is the last item of the session. We take the recommendations from an existing recommendation model depending on the dataset [^15] We chose LLMSeqSim for Amazon Beauty. similarly to Section 4.2, and shuffle them. Then, in the prompt completion we provide the recommendations in the order produced by the model and ensure that the ground truth item is at the top of the list, putting or moving it there as necessary. At prediction time, we ask the fine-tuned model to re-rank the recommendations in the prompt based on what would be the best completion for a prompt session. Similarly to the ranking problem formulation, no post-processing of recommendations is required, since the model chooses between options provided in the prompt.
5. LLM2Sequential: LLM-enhanced Sequential Models
In the third approach, our goal is to leverage the semantically-rich item representations provided by an LLM to enhance an existing sequential recommendation model. Existing recommendation models typically operate on unique identifiers of the items of a domain by learning motifs between the items from the interaction data they are trained with. By equipping the models with meaningful item representations we aim to enable them to learn deeper relationships that can lead to more advanced recommendations.
We reinforce a number of sequential models with LLM capabilities, namely BERT4Rec (Sun et al., 2019), SASRec (Kang and McAuley, 2018), GRU4Rec (Hidasi et al., 2016), and SKNN (Ludewig and Jannach, 2018). Of those, BERT4Rec, SASRec, and GRU4Rec are neural recommendation models that operate on item embeddings. Specifically, these models feature an embedding layer that accommodates embedding representations of items. We supply item embeddings retrieved from an LLM as input to the embedding layer of these models. On the other hand, SKNN is a neighborhood-based sequential model that computes session similarity based on the presence of common items included in the sessions. In this work we adapt SKNN to consider another notion of session similarity grounded on the embedding similarity of items comprising a session.
Besides the consideration of different sequential models, we explore different configurations as done for LLMSeqSim described in Section 3:
  • we use alternative LLM embedding models, i.e., those from OpenAI and Google;
  • we explore different dimensionality reduction methods.
Following the above-mentioned overview of our work on the enhancement of sequential models with LLM capabilities, we organize the description of this approach in two parts. First, we summarize each of the sequential models we consider (Section 5.1). Then, we elaborate how we incorporate LLM embeddings to the workings of the models (Section 5.2).
5.1. Baseline sequential models
We provide a concise description of the baseline sequential models we target, namely SKNN, GRU4Rec, SASRec, and BERT4Rec. Figure 7 depicts the main architectural components of the neural models. We note all the details of our neural model implementations, including any deviations from the original model implementations, in the online material. [^16] https://github.com/dh-r/LLM-Sequential-Recommendation/blob/main/online_material.md
SKNN (Session-based kNN) finds items to recommend by considering sessions similar to the input session(Ludewig and Jannach, 2018). Given a similarity function $\mathrm{sim}(S_{1},S_{2})$ between two sessions $S_{1}$ and $S_{2}$, the model considers the top-$k$ most similar neighboring sessions $N_{S}$ of a session $S$ and computes the recommendation score of an item $i$ according to the following equation, where $\bm{1}(\cdot)$ represents the indicator function.
(1) $\mathrm{score}(i,S)=\sum_{S^{\prime}\in N_{S}}\mathrm{sim}(S^{\prime},S)\cdot\bm{1}(i\in S^{\prime})$
The scores from Equation 1 can be normalized to produce the probability distribution $P(s_{n+1}|S)$. Though the similarity function is intentionally left open, cosine similarity between the sessions’ binary interaction vectors has proven to work well and lead to competitive results in the literature (Ludewig et al., 2021; Kersbergen et al., 2022).
GRU4Rec employs a recurrent neural network based on GRU (Cho et al., 2014). Its overall architecture consists of an item embedding layer, a GRU layer, and a prediction network. GRU4Rec is trained by sequentially processing sessions, where at each timestep it uses the current item in the session and a hidden state computed from previous sessions to predict the item at the next timestep. To elaborate, the item embedding layer consumes the item ID $s_{t}$ at the current timestep $t$ and returns its associated embedding $e_{s_{t}}$ from the item embedding matrix $E$. This item embedding matrix has dimensions $|I|\times e$, where $|I|$ is the number of items and $e$ is the embedding dimension. Then the embedding $e_{s_{t}}$ is fed to a GRU layer. A GRU layer is stateful in that it maintains a hidden state $h_{t}$ while processing a session. This allows the model to use information from the previously seen items $s_{0},s_{1},\ldots,s_{t-1}$ next to item $s_{t}$ to predict $s_{t+1}$. Hence, the GRU layer computes $h_{t}$ based on its previous hidden state $h_{t-1}$ and $s_{t}$. Finally, the model’s output is a score $R_{t}$ for each item, which represents the model’s confidence for each item that it will be the item in the session on timestep $s_{t+1}$. Using $R_{t}$ we can compute next-item recommendations by taking the items with the top-$k$ scores on the last timestep.
In the original paper (Hidasi et al., 2016) the authors experimented with an embedding layer but dropped it and favored a one-hot item encoding because of slightly better performance. We kept the embedding layer to be able to draw a meaningful comparison with other models using LLM embeddings. We elaborate the model’s implementation details in the online material.
SASRec is the first transformer architecture (Vaswani et al., 2017) designed for the sequential recommendation task. It consists of an embedding layer, one or multiple transformer layers, and a prediction network. The transformer layers are based on the encoder stack introduced in (Vaswani et al., 2017). In short, each of the $L$ transformer layers consists of a self-attention (SA) and a feed-forward network (FFN) module. The model is trained by predicting the identity of $s_{t+1}$ for each timestep $t$ like GRU4Rec. However, SASRec computes $R$ for all timesteps $t\in[1,N]$ at once. $R$ is a $N\times I$ matrix, where $R_{t,i}$ denotes the model’s confidence that item $i$ will be the item at position $t+1$.
Given a sequence $S$, either from the embedding layer or a preceding transformer layer, the self-attention module takes a weighted average of all the embeddings in the sequence, for each timestep in the sequence. Given the output sequence of embeddings $S^{L}$ of the last transformer layer, the prediction network of SASRec projects this sequence into scores over all the items. For next-item prediction, we therefore use $R_{N}$, which denotes the vector of the last item $N$ in a session and contains the scores for all items to be the item succeeding the last item in the session.
BERT4Rec is a state-of-the-art neural recommendation model, which employs the transformer architecture (Vaswani et al., 2017) of BERT (Devlin et al., 2019). BERT’s transformer architecture consists of an embedding layer, a stack of transformer layers, and a prediction network. Furthermore, BERT features a masked language model training protocol, which involves masking items at random positions and letting the model predict their true identity.
Initially, the embedding layer embeds an input sequence of (potentially masked) item IDs into a sequence of embeddings using both the item ID and the item position. Then the transformer encoder layers process the embedding sequence using a multi-head attention module and a feed-forward network shared across all positions. Finally, the projection head projects the embeddings at each masked position to a probability distribution in order to obtain the true identity of the masked item. The projection head reuses the item embeddings of the embedding layer to reduce the model’s size and to avoid overfitting.
Figure 7. Architectural components of GRU4Rec, SASRec, and BERT4Rec.
Figure 8. Neural sequential models enhanced with LLM embeddings.
5.2. LLM-enhanced Sequential Models
In this approach, we incorporate LLM embeddings to two different classes of widely applied sequential models: neural sequential models represented by BERT4Rec, SASRec, and GRU4Rec, and neighborhood-based ones represented by SKNN.
As we depict in Figure 8, to allow the neural sequential models to leverage the rich information encoded in LLMs, we employ LLM embeddings for the initialization of the neural models’ item embeddings located in the embedding layer. In order to align the embedding dimension of the LLM embeddings (e.g., 1536) with the configured dimension of the neural models’ embedding layer (e.g., 64), we employ and assess different dimensionality reduction methods. Finally, we train the enhanced model in the same manner as our baseline neural sequential model. The LLM embeddings and dimensionality reduction methods used are described in Section 3.
To enable neighborhood-based models, and SKNN in particular, to use LLM embeddings, we modified the model such that it can operate on the item embeddings. Specifically, we compute a session embedding for each training session. To compute a score for predicting item $i$ of a prompt session $S$ we take the session embedding $e_{S}$ of session $S$ and then consider the similarity of $e_{S}$ with the session embedding $e_{S^{\prime}}$ of each session $S^{\prime}$ of the top-$k$ nearest neighboring sessions $N_{S}$. We then arrive at Equation 2, which differs from Equation 1 in that the latter computes the similarity between two sessions based on the co-occurence of items in these two sessions, while here we derive the similarity from the session embeddings. We experiment with a number of aggregation methods for producing a session embedding from the individual item embeddings of a session. These aggregation methods are the same as the ones used to produce a session embedding from individual item embeddings in LLMSeqSim (Section 3). In our experiments, we also allow the model to use one embedding combination strategy for the training sessions and a different one for the prompt session.
(2) $\mathrm{score}(i,S)=\sum_{S^{\prime}\in N_{S}}\mathrm{sim}(e_{S^{\prime}},e_{S})\cdot\bm{1}(i\in S^{\prime})$
6. Hybrids
The three top-level approaches that we elaborated in Sections 3 (LLMSeqSim), 4 (LLMSeqPrompt), and 5 (LLM2Sequential) have unique characteristics.
LLMSeqSim:
relies on the semantic similarity provided by an LLM between items and compiles a session embedding with desirable properties across items of a session for producing pertinent next-item recommendations.
LLMSeqPrompt:
compiles fundamental and domain-specific knowledge into a single model potentially enabling recommendations that share information elements from both origins.
LLM2Sequential:
is based on incorporating LLM embeddings to the sequential models. This approach introduces meaningful item representations to the training process of sequential models, aiming to boost their learning efficacy towards advanced recommendations.
In this section, we combine the unique characteristics of LLMSeqSim and LLM2Sequential to create stronger hybrids. The hybrids capitalize on the notion of item popularity. Item popularity enables segregating the item space between widely selected items, for which a lot of information is available, and niche items with scant interactions. On the models’ side, we have candidate models that —as we will see— perform very well on popular items (e.g., LLM2Sequential), while other models are agnostic to item popularity and base their recommendations on other signals (e.g., LLMSeqSim). Thus, we can use the popularity property of items to select an appropriate model based on a cutoff popularity threshold. This threshold can be applied either directly on the recommendations of a model (Section 6.2) or alternatively on the last item of a prompt session under the premise that the last item is the driving factor for the next item that will follow (Section 6.1). In this paper, we motivate and research the potential of two concrete hybrids [^17] We also experimented with a hybrid based on embedding model variations, which did not perform better than the individual variations. We share the details in the online material. based on the aforementioned models and the notion of item popularity. We describe these hybrids below.
6.1. Embedding & Sequential Hybrid
Based on the intuition that similarity-based and content-based approaches can lead to better coverage and novelty (Vargas and Castells, 2011; Jannach et al., 2015b; Mendoza and Torres, 2020), while many state-of-the-art recommendation models maintain a bias to popular items, we devise a popularity-based hybrid based on two models. The semantic item recommendation model using LLM embeddings (LLMSeqSim) focuses solely at the semantic representation of items and is oblivious to the notion of item popularity, while LLM2Sequential models are generally good at recommending popular items. In support of this motivation, Figure 9 shows the performance of different models with respect to item popularity. Indeed LLM2BERT4Rec and SKNNEmb exhibit higher hit rates when they recommend for prompt sessions whose last item is popular. In contrast, the hit rate of LLMSeqSim is stable across the popularity of the last item in the prompt sessions.
In practice, we can create a hybrid where LLMSeqSim is used to recommend when encountering unpopular items and LLM2BERT4Rec or SKNNEmb are used to recommend when dealing with popular items at the end of a prompt session. Specifically, we apply a cutoff point on the popularity of the last item of a prompt session and use LLMSeqSim to provide recommendations for that prompt session if the popularity of the last item falls under the cutoff point or one of the two aforementioned sequential models otherwise. Our intuition for using the last item as a driver for the prediction is founded on its impact on correct recommendations according to our analysis in Figure 10, which shows the performance of a model if we treat different positions in a prompt session as last. The figure clearly shows that the item in the last position bears the strongest impact on the performance of recommendations. We apply hyperparameter search on the training set to identify the best cutoff point for popularity considering the quantiles of item popularity per dataset.
Figure 9. Performance per last item popularity.
Figure 10. Performance per item position in session.
6.2. Embedding & Popularity Hybrid
The LLM-embedding based recommendation model (LLMSeqSim) leverages a semantically rich representation of items, but completely lacks desirable notions of recommendations, such as the popularity and diversity of items that traditional recommendation models learn to embrace during their training. Therefore, we create a hybrid based on LLMSeqSim that is also aware of popularity.
The hybrid considers popularity by enforcing a threshold on it such that LLMSeqSim does not recommend obscure items. The threshold could be made relative to the number of options that the model has picked for recommendations in order to include highly relevant recommendations of low popularity. But we did not further explore this alternative. As with the previous hybrid (Section 6.1), we apply hyperparameter search on the training set to determine the optimal popularity cutoff point. We also apply a threshold on diversity in order to enforce the semantic diversity of items in the top-$k$ list of recommendations quantified by the items’ embedding similarity. Specifically, we remove any item whose pair-wise embedding similarity with any other item in the recommendation list exceeds the threshold. We tried both diversity alone and combined with popularity, but it did not produce any promising results. Therefore, we proceeded with popularity only. We share all results in the online material.
7. Experimental Evaluation
In this section we describe our experimental setup (Section 7.1), the results of our empirical evaluation (Section 7.2), and observations from hyperparameter tuning (Section 7.3). We publicly share the code and data of our experiments to ensure reproducibility.
7.1. Experimental setup
The experimental setup consists of the datasets and their pre-processing (Section 7.1.1), the metrics used to evaluate experimental results (Section 7.1.2), the models that were included in the experiments (Section 7.1.3), the hyperparameter tuning process for executing models with their best configurations (Section 7.1.4), and precise information regarding which models were executed on which datasets, including exceptions (Section 7.1.5).
7.1.1. Datasets and Data Splitting
We use the public Amazon Beauty (He and McAuley, 2016) dataset [^18] https://cseweb.ucsd.edu/~jmcauley/datasets/amazon/links.html, Section Per-category file., a novel, real-world proprietary dataset from Delivery Hero, and the public Steam dataset [^19] https://cseweb.ucsd.edu/~jmcauley/datasets.html#steam_data; Version 2: Review Data and Item Metadata. from the gaming domain. The Beauty dataset contains product reviews and ratings from Amazon. In line with prior research (Anelli et al., 2022), we pre-processed the dataset so that there are at least five interactions per user and item ($\mathop{p-\mathrm{core}}=5$). The Delivery Hero dataset contains anonymous QCommerce sessions for dark store and local shop orders. [^20] QCommerce is a segment of e-Commerce focusing on fast delivery times on the last mile. To better simulate a real-world setting, we did not pre-process this dataset, except that we removed sessions with only one interaction. In the Steam dataset we applied $\mathop{p-\mathrm{core}}=5$ pre-processing and removed interactions that contain items with insufficient metadata (title, genres, and tags). Dataset statistics are provided in Table 2.
We selected a diverse set of datasets to ensure that our findings are not limited to datasets with particular characteristics. The datasets are of different nature, originate from different domains, and have different features. In the left part of Figure 11, we visualize the popularity of items in the sessions. After counting the number of occurrences of each item, we rank them in descending order (the most popular being the first) and then plot the log of the frequency on the log of the rank. The Delivery Hero dataset has a tail of items that appear 1 to 4 times, as it is not $\mathop{p-\mathrm{core}}=5$. In the right part of the figure, we show the log of the frequency of the sessions’ length. The Delivery Hero dataset has shorter sessions than the Steam dataset, and fewer longer sessions (those with more than 40 items) than either of the other two datasets.
Figure 12 sheds more light on the differences between the datasets by focusing on the semantic affinity of the items in each session (i.e., on intra-session diversity). To find the semantic affinity of a session, we obtained the OpenAI embeddings of its items. These being vectors, we calculated the pairwise Euclidean distances of the embeddings of each session, and from the pairwise Euclidean distances we then calculated their mean. Greater distances (and their mean) correspond to smaller semantic affinity, or more diverse session items. In this sense, the mean of the pairwise Euclidean distances is a measure of diversity of a session. In the histograms on the left part of Figure 12, we see that the Delivery Hero dataset has higher and less dispersed mean distances. This is also shown in the corresponding boxplots on the right. Less diverse sessions mean that the recommendation problem has to deal with more similar items and thus may be easier to tackle by recommendation models.
Dataset # sessions # items # classes # interactions Avg. length Density
Beauty $\mathop{p-\mathrm{core}}=5$ 22,363 12,101 220 198,502 8.9 0.073%
Delivery Hero 258,710 38,246 NA 1,474,658 5.7 0.015%
Steam $\mathop{p-\mathrm{core}}=5$ 279,290 11,784 700 3,456,395 12.4 0.105%
Table 2. Dataset statistics
Figure 11. Distribution of items ranked by popularity (left) and histogram of session length (right) for the datasets.
Figure 12. Distribution of mean embedding distances for the datasets.
To apply LDA on the data points as explained in Section 3.2, it is required to provide a set of input classes. Recall that the purpose of LDA is to project the data associated with $k$ classes to $k-1$ dimensions so that the projected data best fit those classes. The Beauty and Steam [^21] We use the genres. datasets contain multiple categorical keywords per item. We sort these keywords and concatenate them using underscore as a separator to obtain unique classes per item. The Steam dataset also contains tags, which are provided by users. We process these in the same manner as the genre keywords. The Delivery Hero dataset does not readily include metadata that we can use as classes, so we do not apply LDA to it.
To create a training and test set in a sound way, we first split a dataset containing sessions temporally such that all test sessions succeed train sessions in time. Then in the test set, we adopt the leave-one-out approach as in (Kang and McAuley, 2018; Sun et al., 2019), where all but the last interaction of each session represent the prompt (i.e., the ongoing session), and the last interaction serves as the ground truth.
7.1.2. Metrics
We use the standard ranking accuracy metrics NDCG, MRR, and HitRate at the usual cutoff lengths of 10 and 20. Furthermore, we consider the following beyond-accuracy metrics to obtain a more comprehensive picture of the performance of the different algorithms: catalog coverage, serendipity, and novelty. Catalog coverage represents the fraction of catalog items that appeared in at least one top-$n$ recommendation list of the users in the test set (Jannach et al., 2015b). Serendipity measures the ratio of correct recommendations per user that are not recommended by a popularity baseline (Ge et al., 2010). Novelty computes the negative log of the relative item popularity, or self-information (Zhou et al., 2010).
7.1.3. Models
We include both session-based algorithms of different families, GRU4Rec (Hidasi et al., 2016), and SKNN (Jannach and Ludewig, 2017), as well as two state-of-the-art sequential models, BERT4Rec (Sun et al., 2019) and SASRec (Kang and McAuley, 2018). We tested all variants of the SKNN nearest-neighbor method proposed in (Ludewig and Jannach, 2018) and report the results in the online material. In addition, we include the four LLM-based approaches proposed in Sections 34,  5, and 6. Finally, we include a popularity-based baseline (MostPopular) in the experiments.
7.1.4. Hyperparameter Tuning
We systematically tuned all models (except LLMSeqPrompt) on three validation folds with the Tree Parzen Estimator (TPE) sampler (Bergstra et al., 2011), and the total NDCG@20 across the folds as the optimization goal. We chose to optimize for 20 recommendations to provide enough room for the models to manifest their performance differences, although top-10 and top-20 metrics are usually closely related. In the beginning, we explore 40 random configurations, which helps avoid local minima, and then continue by evaluating the suggestions by the TPE sampler. We let the hyperparameter search run for 72 hours, but started multiple hyperparameter search experiments in parallel for each model (two for vanilla neural models and four for LLM2Sequential models) in order to minimize the duration of experiments. We also allow for early stopping of the hyperparameter search if the optimization objective has not been improved for 100 trials. Finally, we prune a trial if its objective value (i.e., NDCG@20) belongs to the bottom 20% of all tested trials up to that point, allowing us to accelerate the hyperparameter search process. We allow pruning only after ten different configurations of a model have been tried.
Overall, our chosen hyperparameter tuning period with early stopping and pruning enabled allows a wide search of the hyperparameter search space for the various models. Table 3 contains the number of different configurations tried per model. Note that the LLM2Sequential neural models manifest a considerably higher number of trials than the respective vanilla neural models in order to account for the two additional hyperparameters they include, i.e. the embedding models and the dimensionality reduction methods, which result in a search space that is eight times larger.
The folds themselves are created by splitting the training data into validation-training and validation-test sets. We do this in a temporal fashion (same as for the train-test split) by splitting the training sessions into four bins, where each bin contains 25% of the training sessions in chronological order. The first fold uses the first bin as the validation-training data, and the second bin as the validation-test data. Then the second fold uses the first two bins as the validation-training data, and the third bin as the validation-test data. Similarly, the third fold uses the first three bins as the validation-training data, and the fourth bin as the validation-test data. By using less data in the first few folds (in comparison to cross-validation), we can more quickly evaluate a configuration as well as preserve the temporal split in our hyperparameter search process.
For the fine-tuning step of LLMSeqPrompt variants we stick to the default parameters, such as learning rate multiplier and number of epochs, recommended by the LLM providers. In the prediction step where we query a fine-tuned model to provide recommendations, we vary the temperature of the model. This hyperparameter, which in GPT takes values between 0 and 2, controls the randomness of the model’s output. With a lower temperature the model produces more deterministic completions, while higher temperature makes the completions more random. Note that we treat the different LLMs used in the fine-tuning, i.e., OpenAI ada, OpenAI GPT, and Google PaLM, as hyperparameters and report the best performance per task variant. Thus, the number of trials for an LLMSeqPrompt variant in Table 3 represents the temperature configurations tried for all LLMs. The examined hyperparameter ranges and optimal values for each dataset are reported in the online material.
Finally, the hyperparameter tuning of the hybrids involves the following main parameters: embedding source, embedding dimensionality reduction, sequential model (for the LLMSeqSim &Sequential hybrid), and popularity threshold value.
7.1.5. Model Executions on Datasets
In principle, we ran all models on all datasets except for fine-tuning models due to budget and time limitations. Specifically, we applied LLMSeqPromptGenList, LLMSeqPromptClass, and LLMSeqPromptRank only on the Beauty dataset. In addition, we only fine-tuned OpenAI ada on LLMSeqPromptGenItem because it is no longer available for fine-tuning.
Besides fine-tuning experiments, we executed again all experiments of the previous paper (Harte et al., 2023) and in fact managed to tune better a number of baseline models on the Delivery Hero dataset, of which BERT4Rec and SKNN exhibited high performance gains. The performance of SASRec and GRU4Rec also improved moderately. That said, we did observe a small decrease to the performance of SKNN and BERT4Rec in the Amazon Beauty dataset, while GRU4Rec’s performance saw a small increase. We attribute the slightly lower performance to a potential modest instability caused by the size of the dataset and to the stochasticity of the hyperparameter search process.
Model Sec. # Trials Description
BT DH ST
LLMSeqSim 3 591 620 631 Semantic item recommendation model based on OpenAI or Google embeddings
LLMSeqPromptGenItem 4.1 16 14 4 Fine-tuned model to generate the next item
LLMSeqPromptGenList 4.2 5 - - Fine-tuned model to generate next-item recommendations
LLMSeqPromptClass 4.3 5 - - Fine-tuned model to classify items for next-item recommendations
LLMSeqPromptRank 4.4 5 - - Fine-tuned model to rank next-item recommendations
BERT4Rec 5.1 661 52 57 Bidirectional neural model based on the transformer architecture
SASRec 5.1 413 167 146 Unidirectional neural model based on the transformer architecture
GRU4Rec 5.1 622 168 164 Recurrent neural model based on Gated Recurrent Unit
SKNN 5.1 553 377 291 Neighborhood-based sequential model
LLM2BERT4Rec 5.2 1074 118 95 BERT4Rec initialized with LLM embeddings
LLM2SASRec 5.2 1280 272 226 SASRec initialized with LLM embeddings
LLM2GRU4Rec 5.2 1001 381 256 GRU4Rec initialized with LLM embeddings
SKNNEmb 5.2 767 218 128 SKNN utilizing LLM embeddings to compute session similarity
LLMSeqSim &Sequential 6.1 72 72 72 Hybrid of two models that perform well on popular & unpopular items respectively
LLMSeqSim &Popularity 6.2 60 60 60 Semantic item recommendation model enriched with popularity
Table 3. The recommendation models used in the experiments including references to the section where they are described, the number of trials carried out for each dataset (BeauTy, Delivery Hero, STeam) per model in the hyperparameter tuning process, and a short description of each model.
7.2. Results and Discussion
Tables 45, and 6 depict the results obtained for the Amazon Beauty, Delivery Hero, and Steam datasets on the hidden test set, respectively. The rows in each table are sorted in descending order according to the NDCG values at 20. We discuss the experimental results in terms of accuracy and beyond accuracy metrics in Sections 7.2.1 and 7.2.2, respectively.
7.2.1. Accuracy metrics
The key observation of the experimental results is that LLM embeddings boost recommendation model performance in terms of accuracy metrics across models and datasets. Specifically, sequential models enhanced with LLM embeddings score the best performance across all three datasets demonstrating the superiority of LLM embeddings for the models’ learning and training routines. In particular, LLM2SASRec and LLM2BERT4Rec maintain consistently the best performance and alternate on the top spots marking around 45% average improvements in NDCG@20 over their vanilla counterparts on Beauty, and 9% on the Delivery Hero dataset. On Steam, the performance boost is steady but marginal across all sequential models. One potential reason for that could be the quality of embeddings for this particular dataset. As a gaming dataset, Steam features game titles that are semantically non-discriminative. Although we enriched the title with tags that accompany the games, such as action game and multi-player, these are commonly shared across the games and, therefore, their contribution is ambivalent. Finally, LLM2GRU4Rec achieves a 20% increase in NDCG@20 on Beauty over GRU4Rec, but only marginal improvements in the other two datasets.
Top@10 Top@20
Model nDCG HR MRR CatCov Seren Novel nDCG HR MRR CatCov Seren Novel
LLM2SASRec 0.045 0.083 0.034 0.204 0.081 11.660 0.054 0.118 0.036 0.300 0.112 11.844
LLM2BERT4Rec 0.042 0.080 0.031 0.226 0.076 11.645 0.052 0.118 0.034 0.328 0.111 11.823
LLMSeqSim &Sequential 0.040 0.074 0.029 0.349 0.072 11.729 0.050 0.116 0.032 0.514 0.110 11.892
LLMSeqSim 0.043 0.068 0.036 0.761 0.068 13.755 0.049 0.090 0.037 0.889 0.090 13.796
LLMSeqSim &Popularity 0.042 0.065 0.035 0.464 0.065 13.067 0.049 0.093 0.037 0.496 0.092 13.098
SKNNEmb 0.039 0.070 0.030 0.476 0.068 11.855 0.047 0.101 0.032 0.689 0.096 12.090
V_SKNN 0.036 0.063 0.028 0.399 0.059 11.211 0.043 0.090 0.030 0.613 0.083 11.481
LLMSeqPromptRank 0.037 0.068 0.027 0.744 0.068 13.727 0.042 0.089 0.028 0.887 0.089 13.799
LLM2GRU4Rec 0.034 0.065 0.024 0.252 0.062 11.732 0.042 0.097 0.026 0.365 0.091 11.902
BERT4Rec 0.033 0.064 0.023 0.108 0.061 11.789 0.041 0.096 0.025 0.161 0.089 11.937
LLMSeqPromptGenItem 0.032 0.054 0.025 0.548 0.054 13.140 0.037 0.075 0.027 0.728 0.074 13.385
GRU4Rec 0.026 0.051 0.018 0.103 0.046 11.349 0.033 0.080 0.020 0.156 0.073 11.494
SASRec 0.023 0.047 0.016 0.071 0.043 11.081 0.033 0.085 0.018 0.112 0.078 11.316
LLMSeqPromptGenList 0.029 0.053 0.022 0.703 0.053 13.587 0.033 0.069 0.023 0.831 0.069 13.652
LLMSeqPromptClass 0.006 0.012 0.005 0.016 0.010 11.041 0.008 0.017 0.005 0.016 0.012 11.072
MostPopular 0.005 0.010 0.003 0.001 0.001 9.187 0.006 0.018 0.003 0.002 0.001 9.408
Table 4. Evaluation results for the Amazon Beauty dataset
Top@10 Top@20
Model nDCG HR MRR CatCov Seren Novel nDCG HR MRR CatCov Seren Novel
LLM2BERT4Rec 0.101 0.180 0.077 0.240 0.151 10.832 0.120 0.253 0.082 0.301 0.198 11.028
LLMSeqSim &Sequential 0.099 0.175 0.075 0.276 0.146 10.823 0.117 0.247 0.080 0.354 0.191 11.021
SKNNEmb 0.098 0.166 0.077 0.364 0.147 10.938 0.114 0.228 0.082 0.434 0.184 11.076
SKNN 0.098 0.168 0.077 0.361 0.138 10.547 0.113 0.228 0.081 0.438 0.171 10.742
LLM2SASRec 0.096 0.171 0.073 0.275 0.142 10.782 0.113 0.237 0.078 0.358 0.181 10.977
BERT4Rec 0.095 0.169 0.0073 0.278 0.140 10.904 0.112 0.237 0.077 0.364 0.182 11.089
SASRec 0.087 0.153 0.067 0.153 0.126 10.784 0.103 0.217 0.071 0.197 0.163 10.997
LLM2GRU4Rec 0.087 0.154 0.066 0.244 0.127 10.917 0.103 0.218 0.071 0.314 0.163 11.109
GRU4Rec 0.086 0.153 0.066 0.232 0.126 10.869 0.103 0.217 0.070 0.304 0.164 11.058
LLMSeqPromptGenItem 0.063 0.116 0.047 0.400 0.107 12.048 0.070 0.144 0.049 0.611 0.123 13.788
LLMSeqSim &Popularity 0.051 0.091 0.039 0.270 0.090 13.869 0.060 0.126 0.042 0.369 0.122 13.973
LLMSeqSim 0.040 0.073 0.030 0.684 0.072 16.216 0.047 0.102 0.032 0.818 0.099 16.454
MostPopular 0.024 0.049 0.017 0.000 0.000 7.518 0.032 0.079 0.019 0.001 0.000 7.836
Table 5. Evaluation results for the Delivery Hero dataset
Top@10 Top@20
Model nDCG HR MRR CatCov Seren Novel nDCG HR MRR CatCov Seren Novel
LLM2SASRec 0.068 0.129 0.050 0.258 0.112 9.885 0.085 0.197 0.054 0.338 0.150 10.026
LLM2BERT4Rec 0.068 0.128 0.049 0.266 0.110 9.662 0.084 0.194 0.054 0.345 0.143 9.862
BERT4Rec 0.066 0.126 0.047 0.242 0.106 9.583 0.083 0.194 0.052 0.317 0.141 9.813
LLMSeqSim &Sequential 0.066 0.127 0.048 0.410 0.109 9.691 0.083 0.192 0.053 0.548 0.142 9.834
LLM2GRU4Rec 0.066 0.124 0.048 0.319 0.108 9.907 0.082 0.188 0.052 0.414 0.141 10.079
GRU4Rec 0.065 0.123 0.048 0.389 0.106 9.976 0.081 0.188 0.052 0.482 0.141 10.161
SASRec 0.065 0.123 0.047 0.168 0.104 9.647 0.081 0.187 0.052 0.232 0.138 9.869
SKNNEmb 0.051 0.099 0.037 0.243 0.072 8.789 0.064 0.151 0.041 0.353 0.093 9.053
S_SKNN 0.048 0.097 0.034 0.178 0.066 8.439 0.062 0.150 0.038 0.263 0.084 8.684
LLMSeqPromptGenItem 0.030 0.062 0.021 0.457 0.057 10.595 0.037 0.089 0.023 0.686 0.068 11.695
MostPopular 0.022 0.042 0.016 0.001 0.000 7.638 0.030 0.075 0.018 0.002 0.001 7.951
LLMSeqSim &Popularity 0.024 0.044 0.017 0.328 0.041 11.349 0.030 0.068 0.019 0.483 0.059 11.397
LLMSeqSim 0.023 0.044 0.017 0.656 0.041 11.871 0.029 0.067 0.019 0.786 0.058 11.950
Table 6. Evaluation results for the Steam dataset
Furthermore, SKNNEmb outperforms SKNN on Beauty by 9% at NDCG@20 and 12% at HR@20. However, in the other datasets, SKNNEmb only slightly improves over SKNN. A notable observation is that SKNNEmb takes the third place across all models in the Delivery Hero dataset and exhibits the best MRR@10 metric. This result is in line with previous performance comparisons of sequential recommendation models. In (Ludewig and Jannach, 2018), for example, it was observed that the rather simple Sequential Rules (SR) method can lead to highly competitive results in terms of the MRR for some datasets, but is not always competitive in terms of Hit Rate. A strong performance on the MRR measure means that an algorithm is able to frequently position one positive item in the recommendation lists, e.g., due to a popularity bias of the algorithm as in the case of the SR method. Overall, while most accuracy measures are typically highly correlated (Anelli et al., 2022), some algorithms often perform particularly well on certain metrics, depending on specific dataset characteristics. In this respect, our results thus exhibit phenomena that are similar to those reported in the literature.
The LLMSeqSim &Popularity hybrid closely trailed the top models with a small uplift in HR@20 compared to LLMSeqSim on Beauty. However, for the other two datasets, it was not competitive, trailing even the popularity baseline on Steam. That said, besides the tie in NDCG@20 on Beauty, the LLMSeqSim &Popularity hybrid was consistently better than LLMSeqSim on the other datasets and even achieved a 27% increase on the Delivery Hero dataset, showing that the consideration of popularity can lead to more attractive recommendations. Similar to the LLMSeqSim &Popularity hybrid, LLMSeqSim only exhibited high performance on the Beauty dataset. In fact, LLMSeqSim topped the MRR metric, and the LLMSeqSim &Popularity hybrid almost matched the performance. Finally, our expectation of high performance on the Steam dataset due to the even higher homogeneity of items in a session than the Beauty dataset according to Figure 12, did not materialize.
Of the fine-tuned models, LLMSeqPromptGenItem is the most noteworthy since it surpasses the performance of GRU4Rec and SASRec on Beauty, while LLMSeqPromptGenList matches the performance of GRU4Rec and SASRec on the same dataset. Note that both models have been fine-tuned on OpenAI GPT in a single epoch, which was enough for the validation loss to converge. We highlight a number of notable aspects of the fine-tuned models in Section 7.3. In the remaining datasets, LLMSeqPromptGenItem is by far inferior to mainstream recommendation models, but also markedly better than simplistic baselines like MostPopular. Interestingly, the implicit ranking of recommendations in LLMSeqPromptGenItem, where we repeatedly ask the model to provide single-item recommendations until we have top-$k$, was more effective than letting LLMSeqPromptGenList return a ranked list of top-$k$ recommendations. This indication is in line with the performance of LLMSeqPromptRank, which although positioned seemingly high on the performance board, it actually performed worse than LLMSeqSim, which produced the candidate recommendations that LLMSeqPromptRank reranked. Finally, LLMSeqPromptClass performed almost on par with the popularity baseline, showing little to no potential. On the other hand, we barely experimented with the number of clusters, which is a crucial performance factor for this task specification.
7.2.2. Beyond accuracy metrics
We make the following observations for coverage, serendipity, and novelty. The LLMSeqSim model consistently leads to the best catalog coverage and novelty. This is not too surprising, given the nature of the approach, which is solely based on embedding similarities. Unlike other methods that use collaborative signals, i.e., past user-item interactions, the general popularity of an item in terms of the amount of observed past interactions does not play a role in LLMSeqSim, neither directly nor implicitly. Thus, the model has no tendency to concentrate the recommendations on a certain subset of (popular) items. [^22] We recall that the used novelty measure is based on the popularity of the items in the recommendations. Finally, the serendipity results are aligned with the accuracy measures across the datasets. This generally confirms the value of personalizing the recommendations to individual user preferences, compared to recommending mostly popular items to everyone. [^23] We iterate that our serendipity measure counts the fraction of correctly recommended items that would not be recommended by a popularity-based approach.
Interestingly, LLM2SASRec uplifts all beyond-accuracy metrics compared to SASRec, especially catalog coverage and serendipity where it almost doubles catalog coverage on average and contributes an average of 21% improvement on serendipity across datasets. Also notable is the performance of the LLMSeqSim &Sequential hybrid, which is composed of LLMSeqSim and LLM2BERT4Rec, in terms of catalog coverage. The LLMSeqSim &Sequential contributes a 45% average improvement in catalog coverage compared to LLM2BERT4Rec, while sacrificing only a small margin of performance in terms of accuracy metrics. That said, we note that catalog coverage is a coarse metric that counts an item in the coverage of the item space given just a single occurrence of the item across all top-$k$ recommendations on the test set. Still, the improvement is substantial.
7.3. Observations from Hyperparameter Tuning
Given that our proposed approaches embody a variety of novel technical variations, we devote this section to the relationship between parameters and optimal model performance as revealed by the hyperparameter search process for LLMSeqSim (Section 7.3.1), LLM2Sequential (Section 7.3.2), LLMSeqPrompt (Section 7.3.3), and hybrid models (Section 7.3.4). Table 3 depicts the number of hyperparameter search trials for each model and dataset.
7.3.1. Semantic Item Recommendations via LLM Embeddings (LLMSeqSim)
There are four main observations that stand out from the tuning of LLMSeqSim across datasets. First, the use of OpenAI embeddings leads to a clear performance gain compared to embeddings obtained from Google. Second, dimensionality reduction increases accuracy and, third, the application of PCA always results in the best configuration. Finally, while reducing the dimensions of the embeddings always helps, the higher the target number of dimensions is, the higher the accuracy.
7.3.2. LLM-enhanced Sequential Models (LLM2Sequential)
For the sequential models enhanced with LLM embeddings, the choice of LLM embedding model per se only leads to marginal performance differences. In fact, both OpenAI and Google embeddings are found in best model configurations depending on the dataset and model. On the other hand, dimensionality reduction consistently contributes to the best model configurations across datasets. Usually, PCA is the optimal reduction method, but LDA and random projection also appear in certain cases. Finally, neural sequential models benefit from low learning and drop rates, while SKNNEmb leverages different types of decay for training and prompt sessions to optimize its performance. SKNNEmb is agnostic to the choice of similarity type, likely due to the normalization of embedding vectors.
7.3.3. Fine-tuned Models (LLMSeqPrompt)
A fundamental challenge of fine-tuning regards whether the fine-tuned model will comply with the form of the task at hand. All models that we fine-tuned across datasets closely aligned the form of their responses with the completions provided to them during the fine-tuning process according to the task specifications that we described in Section 4.
We iterate that we applied all four LLMSeqPrompt fine-tuning task variants on Beauty and only LLMSeqPromptGenItem on the Delivery Hero and Steam datasets due to budget and time limitations. Section 7.1 includes all information regarding the fine-tuning experiments. In all setups where both GPT and PaLM are fine-tuned, i.e., for all variants in the Beauty, Delivery Hero, and Steam datasets, GPT outperforms PaLM demonstrating an average uplift of 78% in terms of NDCG@20. We include the corresponding results in the online material. For reference, these observations appear to contradict the Chain of Thought Hub (Fu et al., 2023), which evaluates the reasoning performance of LLMs, where PaLM 2 exhibits noticeably better performance in a variety of reasoning tasks compared to GPT 3.5 turbo. However, PaLM 2 refers to the unicorn version, while we used the bison one, which is smaller than unicorn. On the other hand, GPT 3.5 turbo performs substantially better than PaLM bison in the LMSys Chatbot Arena leaderboard, [^24] https://leaderboard.lmsys.org which evaluates the user dialogue with an LLM based on user votes.
In addition, the LLMSeqPromptGenItem task has been used to fine-tune both ada and GPT 3.5 turbo OpenAI models on the Beauty and Delivery Hero datasets. The results are conflicting. On the one hand, we observe a notable performance increase of 23% in terms of NDCG@20 in favor of GPT on Beauty, but, surprisingly, on Delivery Hero ada performed significantly better than GPT for the same task achieving 18% higher NDCG@20. Given that the comparison scope is narrow, we can only conclude that the choice of fine-tune model depends on the dataset and that the evolution of models does not necessarily lead to higher performance.
With respect to the temperature and top_p hyperparameters we observe that in all setups a high temperature results in lower performance as it allows for more creativity at the cost of item hallucinations. For LLMSeqPromptGenItem and LLMSeqPromptGenList a low but non-zero temperature leads to the best performance whereas LLMSeqPromptClass and LLMSeqPromptRank perform best with temperature set to zero. A possible explanation for this is that the latter variants do not require any creativity because the recommendation options are contained in the provided prompt.
Dataset Task Model HR@20 NDCG@20 # Uniq. Reco. Hallucin. Emb. Sim.
[…]
GPT fntd 0.077 0.038 20056 36.11% 0.960
[…]
LLMSeqPromptGenItem PaLM fntd 0.043 0.020 21103 56.23% 0.950
[…]
GPT fntd 0.069 0.033 47966 43.35% 0.953
[…]
Beauty LLMSeqPromptGenList PaLM fntd 0.042 0.020 32643 21.24% 0.969
[…]
GPT fntd 0.111 0.057 55301 13.79% 0.967
[…]
DH LLMSeqPromptGenItem PaLM fntd 0.050 0.029 43698 37.87% 0.961
[…]
GPT fntd 0.059 0.032 5701 0.36% 0.889
[…]
Steam LLMSeqPromptGenItem PaLM fntd 0.025 0.014 382446 58.87% 0.867
Table 7. Comparison between base and fine-tuned models across datasets for LLMSeqPromptGenItem and LLMSeqPromptGenList in terms of HR@20, NDCG@20, number of unique recommendations, percentage of hallucinations, and embedding similarity between hallucinations and closest match.
Finally, to explore the learning capabilities of fine-tuned models we compare their performance to the corresponding base models. We observe that in all setups the fine-tuned models significantly outperform their non fine-tuned counterparts. For LLMSeqPromptGenItem and LLMSeqPromptGenList variants the model generates recommendations that are not part of the prompt. For those, the base model has a much higher rate of hallucinations compared to fine-tuned models as Table 7 shows. Additionally, the hallucinations produced by the fine-tuned models exhibit a higher embedding similarity to the catalog items compared to the base model. These observations indicate that fine-tuned models recall and prioritize concepts that they processed during their fine-tuning compared to the base model, which is not given the same opportunity. Therefore, the base model has to provide recommendations solely based on the information in the prompt.
For variants LLMSeqPromptClass and LLMSeqPromptRank the recommendation options are explicitly part of the context such that the model only has to select and/or re-rank the given item names. Thus, hallucinations are uncommon in those variants. Still, the performance advantage of the fine-tuned models over the base models persists. This indicates that the fine-tuned models learn the underlying classification and ranking tasks better. This is an expected finding that is widely acknowledged by literature and practice (, 2023; Touvron et al., 2023).
All in all, the fine-tuned models perform better for each recommendation task with less hallucinations and closer semantic proximity to the actual options compared to the base models. This finding clearly shows that fine-tuning can enable a model to learn not only a task, but also concepts of a domain to some extent and contradicts existing findings and common belief that fine-tuning is only applicable for learning form rather than facts (Mitra et al., 2023; Kirstain et al., 2022; Ovadia et al., 2024; Lewis et al., 2021; Wang et al., 2023; Kandpal et al., 2023; Luo et al., 2023; Chen et al., 2020; Kirkpatrick et al., 2017; Berglund et al., 2023).
7.3.4. Hybrids
In both hybrids (LLMSeqSim &Sequential, LLMSeqSim &Popularity) the LLMSeqSim model based on the OpenAI embeddings performs better, while for the LLM2Sequential model performance depends on the dataset: the OpenAI embeddings dominate on the Delivery Hero and Steam datasets, and the Google embeddings on the Beauty dataset. With regard to dimensionality reduction, it mainly improves performance compared to the original embedding size.
With respect to the popularity threshold, for LLMSeqSim &Sequential we note that increasing the cutoff point seems to decrease performance. Thus, the hybrid appears to favor LLM2Sequential over LLMSeqSim since a lower threshold results in more recommendations from the former, and less from the latter. In contrast, in LLMSeqSim &Popularity increasing the popularity threshold value appears to increase performance, especially for the Delivery Hero and Steam datasets.
8. Challenges, Limitations, & Conclusions
While large language models offer a variety of new opportunities to build highly effective (sequential) recommender systems, their use, both in industry and academic research, may come with a number of challenges. One central issue in that context is that the most powerful LLMs today are under the control of large organizations and are primarily accessed via APIs. These LLMs are thus black boxes to their users. Moreover, and maybe more importantly, the behavior of these LLMs can change over time and their performance can vary. In an industrial setting, this may represent a major challenge from a quality assurance perspective. In academic settings, the varying behavior of the models can make the exact reproducibility of reported findings almost impossible.
Furthermore, depending on the application use case, an LLM fine-tuning step may be required to achieve satisfactory recommendation performance. In cases where the organization that hosts an LLM provides an API for fine-tuning, this may lead to data protection and privacy issues, as potentially sensitive company-internal data may have to be uploaded for the fine-tuning step. Finally, depending on the particular way an LLM is leveraged to support (sequential) recommendation processes, significant costs may arise with the use of an LLM. Real-world item catalogs can be huge, and fine-tuning may furthermore be required frequently, e.g., in case of use cases with highly dynamic item catalogs. In this context, we recall that for some datasets it may not be sufficient to feed item names into the LLM, as they may carry limited semantic information, as in the case of the Steam dataset. In such situations, potentially available longer item descriptions may be used, but this approach may further add to the cost of the fine-tuning process.
In terms of research limitations, our work so far is based on three datasets from two domains (e-commerce, games). As the datasets used in our research are markedly different in terms of various characteristics (e.g., sparsity, size, available meta-data), we are confident that our findings generalize beyond the conducted experiments. Nonetheless, further experiments are needed to validate the benefits of leveraging LLMs for sequential recommendation in other domains, e.g., in music or news recommendation.
To conclude, Large Language Models and AI assistants like ChatGPT have exhibited disruptive effects in various domains in the past few years. In this work, we have conducted an extensive analysis on the potential benefits of LLMs for the highly relevant class of sequential recommendation problems. Specifically, we have designed, investigated, and systematically assessed different ways of leveraging LLMs in the recommendation process. Our findings show that existing sequential models can particularly benefit from the semantic knowledge encoded in LLMs, leading to substantial accuracy increases compared to previous models. Furthermore, depending on the application, sometimes the LLM-enhanced retrieval of semantically similar items turned out to be highly effective. Our future work includes both the analysis of the proposed methods for additional domains and datasets and the design of alternative ways of building recommendation approaches based on LLMs.

### v2

@@
 The key contribution of LLMs to the sequential recommendation problem is a semantic model of real world concepts and their representation in natural language sequences.
-This consideration is neglected by typical sequential recommendation models (Kang and McAuley, 2018; Sun et al., 2019).
+This consideration is neglected by typical ID-based sequential recommendation models (Kang and McAuley, 2018; Sun et al., 2019).
 The rapidly growing interest for involving LLMs in recommender systems clusters around cross-domain (Zhang et al., 2021a; Hou et al., 2023b), domain-specific (Bao et al., 2023; Zhang and Wang, 2023), and multi-modal (Wu et al., 2022; Wang et al., 2022) recommendations.
@@
 Altogether, the current state of research is summarized in the following surveys (Lin et al., 2023; Wu et al., 2023).
-Given the currently early stage of this research avenue, the prospect for the various recommendation approaches involving LLMs remains largely unexplored.
-In this work, [^2] The present paper substantially extends our previous work on leveraging LLMs for the sequential recommendation problem (Harte et al., 2023).
-We detail our introduced extensions in Section 2.
-we design and research the potential of three orthogonal approaches and hybrids of those that leverage LLMs for sequential recommendations.
+In this paper, we build on the ideas developed in these early works on involving LLMs in recommender systems, and we design and research the potential of three orthogonal approaches and hybrids of those that leverage LLMs for the particular problem setting of sequential recommendations.
 We decompose each approach into its constituent technical aspects, identify or devise alternatives for each aspect, and investigate the impact of choices to the approach’s performance.
@@
 Our work results in the following contributions and insights.
-  • We design and research three orthogonal methods and two hybrid approaches of leveraging LLMs for sequential recommendation.
+[^2] The present paper substantially extends our previous work on leveraging LLMs for the sequential recommendation problem (Harte et al., 2023).
+We detail our introduced extensions in Section 2.
+  • Building on the growing literature regarding the use of LLMs in recommender systems, we design and research three orthogonal methods and two hybrid approaches of leveraging LLMs for sequential recommendation.
 In particular, we investigate the technical aspects of each method, determine alternatives for each, thoroughly tune them, and evaluate their impact.
@@
 Both results support the potential of LLM-based models for datasets with specific characteristics.
+Overall, the main contribution of our work is an in-depth analysis of the potential of various previously-explored and novel ways of leveraging LLMs for sequential recommendation problems.
+Our comprehensive analyses, which involve different datasets, three LLMs, and a set of orthogonal technical approaches reveal that substantial performance improvements can indeed be obtained by involving LLMs for sequential recommendations.
 The rest of the paper is organized as follows.
@@
 The aim is to predict the next interaction of the given sequence.
-Besides the recent sequential recommendation models mentioned in the introduction (Hidasi et al., 2016; Kang and McAuley, 2018; Sun et al., 2019), in earlier works the sequential recommendation problem has been modelled as a Markov Chain (Garcin et al., 2013) or a Markov Decision Process (Shani et al., 2005).
+Besides the recent sequential recommendation models mentioned in the introduction (Hidasi et al., 2016; Kang and McAuley, 2018; Sun et al., 2019), S3-Rec (Zhou et al., 2020) pre-trains a transformer architecture enhanced with self-supervised signals aiming to learn better data representations using the mutual information maximization principle.
+In earlier works the sequential recommendation problem has been modelled as a Markov Chain (Garcin et al., 2013) or a Markov Decision Process (Shani et al., 2005).
 Neighborhood-based approaches, such as SKNN (Jannach and Ludewig, 2017), have also been proposed.
@@
 Finally, we recommend the top-$k$ products from the catalog with the highest embedding similarity to the session embedding.
+Overall, LLMSeqSim shares similarities with the sequential model proposed in (Li et al., 2023c), which also relies on creating item embeddings from textual representations derived from metadata features, although not based on LLMs.
 We identified four technical aspects in this processing pipeline, where different implementation alternatives are possible.
@@
 4. LLMSeqPrompt: Prompt-based Recommendations by a Fine-Tuned LLM
-In this approach, we inject domain knowledge to the collective information that a base LLM incorporates, with the goal of combining fundamental information about the items of the domain encoded in the LLM with domain-specific information conveyed in the fine-tuning process.
-The composition of these two elements of information exposed in the same highly capable substrate of an LLM paves the path towards potentially advanced recommendations.
-To this end, we fine-tune an OpenAI ada [^7] https://platform.openai.com/docs/guides/fine-tuning model, an OpenAI GPT-3.5 turbo [^8] https://platform.openai.com/docs/guides/fine-tuning/what-models-can-be-fine-tuned model and the model text-bison@001999https://cloud.google.com/vertex-ai/docs/generative-ai/models/tune-models, which is based on Google PaLM2.
+In this approach, we inject domain knowledge to the collective information that a base LLM incorporates.
+Our goal is to thereby combine fundamental information about the items of the domain encoded in the LLM with domain-specific information conveyed in the fine-tuning process.
+Related approaches to fine-tune pre-trained LLMs for recommendation problems in different ways were successfully explored in (Bao et al., 2023; Zhang et al., 2023; Wang et al., 2022).
+To this end, we fine-tune an OpenAI ada [^7] https://platform.openai.com/docs/guides/fine-tuning model, an OpenAI GPT-3.5 turbo [^8] https://platform.openai.com/docs/guides/fine-tuning/what-models-can-be-fine-tuned model, and the model text-bison@001999https://cloud.google.com/vertex-ai/docs/generative-ai/models/tune-models, which is based on Google PaLM2.
 GPT succeeded ada as OpenAI’s recommended base model for fine-tuning, while PaLM2 is a base model created by Google.
 While there is general consensus on the overall architecture and training process of a base LLM, differences in architectural choices, its configuration, and the training data mean that LLMs can differ significantly.
-[^10] A discussion of those factors is out of this paper’s scope.
-Thus, the aforementioned selection of models enables us to compare and contrast two fundamentally different LLMs and also assess the performance evolution and qualitative characteristics of OpenAI models, going from ada to GPT 3.5 turbo.
+Thus, the three selected models enable us to compare and contrast two fundamentally different LLMs, and to also assess the performance evolution of the two OpenAI models.
 We fine-tune the models on training samples consisting of a prompt (the input) and a completion (the intended output).
 First, we take the training sessions and format them as prompt-completion pairs according to the fine-tuning task specification.
-We present the precise task specifications in Sections 4.14.24.3, and 4.4.
+[^10] We present the precise task specifications in Sections 4.1 to Section 4.4.
 We also separate a small subset of the training sessions and provide them to the fine-tuning process as validation set.
@@
 In a classification task, in contrast, a fine-tuned model is presented with a candidate pool of item recommendations presented as categories and selects a top-$k$ subset of those similarly to multi-class classification.
-Finally, the ranking task further narrows the freedom of the model in that the set of available recommendations is provided in the prompt, allowing the model to focus on the best ranking of those recommendations.
+Finally, the ranking task further narrows the freedom of the model in that the set of available recommendations is provided in the prompt.
+This allows the model to focus on the best ranking of those recommendations.
 Figure 3. Next-item generation by fine-tuned LLM.
-4.1. Fine-tuning Task: Generate Single Recommended Item
+4.1. Fine-tuning Task: Generate a Single Recommended Item
 In this variant, the prompt is a session, which contains a list of item metadata, e.g., product names, except for the last item, and the completion contains the metadata of the last item in the same session according to the first step in Figure 3.
 At prediction time, following the fine-tuning process, we repeat a prompt $k$ times ($k=20)$ and collect a single recommendation produced by the model per prompt invocation according to the fine-tune task formulation.
-We use the tendency of the model to make the same recommendation in different invocations for the same prompt as a proxy of its confidence for the recommendation, collect and deduplicate the recommendations over the $k$ invocations and rank the recommendations by frequency of appearance.
+We use the tendency of the model to make the same recommendation in different invocations for the same prompt as a proxy of its confidence for the recommendation.
+We then collect and deduplicate the recommendations over the $k$ invocations and rank the recommendations by frequency of appearance.
 The fine-tuned LLM, being a generative model, may also return hallucinated products.
-To remedy that, in the fifth step, we retrieve the embedding of each deduplicated recommendation using either the OpenAI or the Google embeddings API (this embedding might correspond to a hallucinated product), and finally we take the catalog’s product that is closest in terms of embedding similarity using the dot product.
+To remedy that, in the fifth step, we retrieve the embedding of each deduplicated recommendation and finally take the catalog’s product that is closest in terms of embedding similarity using the dot product.
 Repeating the prompt allows us to get a number of different recommendations from the single recommendation returned each time.
@@
 Figure 4. Recommendation list generation by fine-tuned LLM.
-4.2. Fine-tuning Task: Generate List of Item Recommendations
-By design, a recommendation model aims to produce a number of desirable options in order to maximize the probability of satisfying a user’s intent at the time.
-Therefore, we would like to teach an LLM to embrace the role of a typical recommendation model and generate a list of next-item recommendations ranked by order of relevance that is determined by the fine-tuned model.
+4.2. Fine-tuning Task: Generate a List of Item Recommendations
+In this approach, we teach an LLM to embrace the role of a typical recommendation model and generate a list of next-item recommendations ranked by order of relevance.
 To convey this task in the fine-tuning process as we depict in Figure 4, we provide in the prompt a session containing a list of items except the last one.
 In the prompt completion we provide a list of item recommendations that is compiled by a state of the art recommendation model.
-The rationale is to ask a proven recommendation model to provide recommendations for the training sessions and use those as prompt completions for the fine-tune process.
-In fact, we can query more than one recommendation model and use the recommendations of the model that achieves the highest NDCG, or another metric, per training session.
-With enough training sessions, this practice could potentially transfer the knowledge of a powerful ensemble of models to a fine-tuned LLM.
-[^13] We did not apply this step in this work.
-Instead, we used LLMSeqSim in the Beauty dataset due to its high performance.
+The rationale is to ask a proven recommendation model to provide recommendations for the training sessions and use those as prompt completions for the fine-tuning process.
+In our case, we used LLMSeqSim in the Beauty dataset due to its high performance.
+[^13] In theory, we could also query more than one recommendation model and use the recommendations of the model that achieves the highest NDCG, or another metric.
 Finally, we insert the ground truth item at the top of the list, if not already there, and fine-tune an LLM with the same prompts and the corresponding compiled list of recommendations as completion.
@@
 4.3. Fine-tuning Task: Classify Items for Next Item Recommendations
-Asking an LLM to generate responses without explicitly providing the eligible options in the prompt represents the most challenging type of task.
-The reason is that it forces the model to choose from a virtually infinite set of options often leading to hallucinations.
-Therefore, we aim to liberate an LLM from the complexity of generating unbounded responses based on its fine-tuning.
-Instead, we allow it to focus on picking and ranking a list of pertinent item recommendations from an overall pool of items that we pose to the model as classification categories.
-The intuition behind this formulation is that there exists a modest subset of popular and diverse items of a domain that will please most customers avoiding, at the same time, niche items that receive scant attention.
-In this variant, depicted in Figure 5, we state the next item recommendation problem as a multi-class classification task as follows.
-Prior to the fine-tuning process, we apply a clustering algorithm, namely K-means, to group items to clusters based on the embedding distances of their metadata [^14] We use product names in the Amazon Beauty dataset.
-As a result, the clusters can be seen as a set of diverse item groups that are potentially representative of the whole item space.
-We pick a relatively large number of clusters, i.e., 200, to maximize the probability of a good representation of semantically different items by the corresponding clusters.
+In this approach, we explicitly provide the LLM with the eligible options for recommendation, instead of allowing it to ‘freely’ return recommendations.
+The learning goal is thus to select and rank a set of pertinent item recommendations from an overall pool of items that we pose to the model as classification categories.
+The intuition behind this formulation is that there exists a small subset of popular and diverse items in each domain that will please most customers.
+Restricting the set of options this way also helps us to avoid niche items that receive scant attention.
+We state the next item recommendation problem as a multi-class classification task as follows, see also Figure 5.
+Prior to the fine-tuning process, we apply a clustering algorithm, namely K-means, to group items to clusters based on the embedding distances of their metadata.
+[^14] We use product names in the Amazon Beauty dataset.
+The resulting clusters can be seen as a set of diverse item groups that are representative of the whole item space.
+We pick a relatively large number of clusters, i.e., 200, to maximize the probability of a good representation of semantically different items.
 Then we elevate the most popular item of a cluster to represent the group of items in the cluster.
@@
 Finally, for each training session, we take the embeddings of the ground truth item and the items that represent the classification categories and select the top-$k$ semantically closest classification categories for each ground truth item.
-In the prompts used for fine-tuning, together with each input session of items (except the last item), we include the classification categories and in the corresponding completions we include the top-$k$ classification categories that we computed in the previous step.
+The prompts for fine tuning include the input session of items (except) the last item and the classification categories.
+In the corresponding completions, we include the top-$k$ classification categories that we computed in the previous step, see Figure 5 for an example.
 At prediction time, we ask the model to classify a prompt session to top-$k$ categories, which correspond to next item recommendations.
 Note that since the candidate recommendations are included in the prompt, there is no need for an extra step to map the recommendations back to the item space.
-The fine-tuned models respond according to their fine-tuning task specification.
 Figure 6. Rank recommendation slate by fine-tuned LLM.
 4.4. Fine-tuning Task: Rank Next Item Recommendations
-The ranking task enables the fine-tuned model to draw its attention to the ranking of recommendations.
-By fine-tuning a model to rank recommendations, we aim to investigate the efficacy in the ranking task of an LLM adapted to a particular domain compared to the more open fine-tuning task specifications of Sections 4.14.2, and 4.3.
-In this variant, visualized in Figure 6, we fine-tune the models to learn a ranking task where we pass in the prompt a session of items except the last one and a set of recommendations including the ground truth, that is the last item of the session.
-We take the recommendations from an existing recommendation model depending on the dataset [^15] We chose LLMSeqSim for Amazon Beauty.
+In this final variant, visualized in Figure 6, we fine-tune the models to learn a ranking task.
+Specifically, we create a prompt that consists of (a) a session of items except the last one, and (b) a set of recommendations including the ground truth, that is the last item of the session.
+We take the set of recommendations from an existing recommendation model depending on the dataset [^15] We chose LLMSeqSim for Amazon Beauty.
 similarly to Section 4.2, and shuffle them.
@@
 At prediction time, we ask the fine-tuned model to re-rank the recommendations in the prompt based on what would be the best completion for a prompt session.
-Similarly to the ranking problem formulation, no post-processing of recommendations is required, since the model chooses between options provided in the prompt.
 5. LLM2Sequential: LLM-enhanced Sequential Models
@@
 Then, we elaborate how we incorporate LLM embeddings to the workings of the models (Section 5.2).
+Figure 7. Architectural components of GRU4Rec, SASRec, and BERT4Rec.
 5.1. Baseline sequential models
@@
 The projection head reuses the item embeddings of the embedding layer to reduce the model’s size and to avoid overfitting.
-Figure 7. Architectural components of GRU4Rec, SASRec, and BERT4Rec.
 Figure 8. Neural sequential models enhanced with LLM embeddings.
@@
 In this approach, we incorporate LLM embeddings to two different classes of widely applied sequential models: neural sequential models represented by BERT4Rec, SASRec, and GRU4Rec, and neighborhood-based ones represented by SKNN.
+Neural Models
 As we depict in Figure 8, to allow the neural sequential models to leverage the rich information encoded in LLMs, we employ LLM embeddings for the initialization of the neural models’ item embeddings located in the embedding layer.
@@
 The LLM embeddings and dimensionality reduction methods used are described in Section 3.
+In terms of technical aspects, we note that the item embeddings in our approach are trainable and they are computed based on item metadata of the domain.
+For the Beauty and the Delivery Hero datasets, the metadata used to compute the embeddings are the names of the products.
+In the Steam dataset, the items of the domain are games.
+Typically, the names of the games are not closely linked with the concept of a game.
+Thus, we concatenate the names of the games with tags that accompany and characterize each game.
+The IDs of the items are not included to the input of the embedding computation.
+In the overall process, the collaborative information thus enters the LLM2Sequential model as a sequence of IDs representing the items of the domain that appear in the sequence.
+In the embedding layer, the sequence of IDs becomes a sequence of embeddings in the following manner.
+First, the item embedding of each item in the sequence is retrieved from the embedding matrix based on the ID of the item.
+The item embedding is the embedding previously received by an LLM based on metadata of the item provided as input to the LLM.
+Second, each item embedding is summed with the positional embedding of the item’s position in the sequence to form the final embedding representation.
+Then, the embeddings are transformed while passing through layers of the model in the feed-forward pass where the collaborative and textual information are fused in the multilayer perceptron of the feed-forward network.
+Finally, through backpropagation of the gradients that minimize the cost, the weights of each layer are updated leading to the update of the embeddings in the embedding layer as well.
+Neighborhood-based Models
 To enable neighborhood-based models, and SKNN in particular, to use LLM embeddings, we modified the model such that it can operate on the item embeddings.
@@
 7.1. Experimental setup
-The experimental setup consists of the datasets and their pre-processing (Section 7.1.1), the metrics used to evaluate experimental results (Section 7.1.2), the models that were included in the experiments (Section 7.1.3), the hyperparameter tuning process for executing models with their best configurations (Section 7.1.4), and precise information regarding which models were executed on which datasets, including exceptions (Section 7.1.5).
+The experimental setup consists of the datasets and their pre-processing (Section 7.1.1), the metrics used to evaluate experimental results (Section 7.1.2), the models that were included in the experiments (Section 7.1.3), the hyperparameter tuning process for executing models with their best configurations (Section 7.1.4), the experimental setup of S3-Rec (Section 7.1.5), TALLRec (Section 7.1.6), and LlamaSeqPromptGenItem (Section 7.1.7) as well as precise information regarding which models were executed on which datasets, including exceptions (Section 7.1.8).
 7.1.1. Datasets and Data Splitting
-We use the public Amazon Beauty (He and McAuley, 2016) dataset [^18] https://cseweb.ucsd.edu/~jmcauley/datasets/amazon/links.html, Section Per-category file., a novel, real-world proprietary dataset from Delivery Hero, and the public Steam dataset [^19] https://cseweb.ucsd.edu/~jmcauley/datasets.html#steam_data; Version 2: Review Data and Item Metadata.
-from the gaming domain.
-The Beauty dataset contains product reviews and ratings from Amazon.
-In line with prior research (Anelli et al., 2022), we pre-processed the dataset so that there are at least five interactions per user and item ($\mathop{p-\mathrm{core}}=5$).
-The Delivery Hero dataset contains anonymous QCommerce sessions for dark store and local shop orders.
-[^20] QCommerce is a segment of e-Commerce focusing on fast delivery times on the last mile.
-To better simulate a real-world setting, we did not pre-process this dataset, except that we removed sessions with only one interaction.
-In the Steam dataset we applied $\mathop{p-\mathrm{core}}=5$ pre-processing and removed interactions that contain items with insufficient metadata (title, genres, and tags).
-Dataset statistics are provided in Table 2.
-We selected a diverse set of datasets to ensure that our findings are not limited to datasets with particular characteristics.
-The datasets are of different nature, originate from different domains, and have different features.
-In the left part of Figure 11, we visualize the popularity of items in the sessions.
-After counting the number of occurrences of each item, we rank them in descending order (the most popular being the first) and then plot the log of the frequency on the log of the rank.
-The Delivery Hero dataset has a tail of items that appear 1 to 4 times, as it is not $\mathop{p-\mathrm{core}}=5$.
-In the right part of the figure, we show the log of the frequency of the sessions’ length.
-The Delivery Hero dataset has shorter sessions than the Steam dataset, and fewer longer sessions (those with more than 40 items) than either of the other two datasets.
+[…]
+[…]
 Figure 12 sheds more light on the differences between the datasets by focusing on the semantic affinity of the items in each session (i.e., on intra-session diversity).
@@
 Dataset # sessions # items # classes # interactions Avg. length Density
-Beauty $\mathop{p-\mathrm{core}}=5$ 22,363 12,101 220 198,502 8.9 0.073%
+[…]
 Delivery Hero 258,710 38,246 NA 1,474,658 5.7 0.015%
-Steam $\mathop{p-\mathrm{core}}=5$ 279,290 11,784 700 3,456,395 12.4 0.105%
+[…]
 Table 2. Dataset statistics
@@
 7.1.3. Models
-We include both session-based algorithms of different families, GRU4Rec (Hidasi et al., 2016), and SKNN (Jannach and Ludewig, 2017), as well as two state-of-the-art sequential models, BERT4Rec (Sun et al., 2019) and SASRec (Kang and McAuley, 2018).
-We tested all variants of the SKNN nearest-neighbor method proposed in (Ludewig and Jannach, 2018) and report the results in the online material.
-In addition, we include the four LLM-based approaches proposed in Sections 34,  5, and 6.
+We include both session-based algorithms of different families, GRU4Rec (Hidasi et al., 2016), and SKNN (Jannach and Ludewig, 2017), as well as three state-of-the-art sequential models, BERT4Rec (Sun et al., 2019), SASRec (Kang and McAuley, 2018), and S3-Rec (Zhou et al., 2020).
+Notably, S3-Rec introduces a custom pre-training and fine-tuning method that allows a meaningful comparison to our proposed LLM-based approaches, especially with respect to the LLM embedding initialization of neural models (Section 5) and finetuning of LLMs (Section 4).
+We also provide a comparison with a fine-tuned open-source Llama (Touvron et al., 2023) LLM.
+In addition, we evaluate TALLRec (Bao et al., 2023), a state of the art LLM-based approach for sequential recommendations that structures the recommendations as instructions and tunes an LLM through an instruction-tuning process.
+Both Llama and TALLRec are closely related to our proposed LLM fine-tuning approaches (Section 4).
+Additionally, we initialize neural models with embeddings retrieved from a pre-trained BERT (Devlin et al., 2019) model to compare against our proposed initialization approach using embeddings retrieved from OpenAI GPT and Google PaLM (Section 5).
+We refer to those models as BERT2BERT4Rec, BERT2SASRec, and BERT2GRU4Rec.
+Furthermore, we tested all variants of the SKNN nearest-neighbor method proposed in (Ludewig and Jannach, 2018) and report the results of the best SKNN variant in this paper.
+The complete results can be found in the online material.
+Moreover, we include the four LLM-based approaches proposed in Sections 34,  5, and 6.
 Finally, we include a popularity-based baseline (MostPopular) in the experiments.
+[^22] The consideration of additional traditional baselines and most recently published sequential models is left for future work.
+Model Sec. # Trials Description
+BT DH ST
+LLMSeqSim 3 591 620 631 Semantic item recommendation model based on OpenAI or Google embeddings
+LLMSeqPromptGenItem 4.1 16 14 4 Fine-tuned model to generate the next item
+LLMSeqPromptGenList 4.2 5 - - Fine-tuned model to generate next-item recommendations
+LLMSeqPromptClass 4.3 5 - - Fine-tuned model to classify items for next-item recommendations
+LLMSeqPromptRank 4.4 5 - - Fine-tuned model to rank next-item recommendations
+BERT4Rec 5.1 661 52 57 Bidirectional neural model based on the transformer architecture
+SASRec 5.1 413 167 146 Unidirectional neural model based on the transformer architecture
+GRU4Rec 5.1 622 168 164 Recurrent neural model based on Gated Recurrent Unit
+SKNN 5.1 553 377 291 Neighborhood-based sequential model
+LLM2BERT4Rec 5.2 1074 118 95 BERT4Rec initialized with LLM embeddings
+LLM2SASRec 5.2 1280 272 226 SASRec initialized with LLM embeddings
+LLM2GRU4Rec 5.2 1001 381 256 GRU4Rec initialized with LLM embeddings
+SKNNEmb 5.2 767 218 128 SKNN utilizing LLM embeddings to compute session similarity
+LLMSeqSim &Sequential 6.1 72 72 72 Hybrid of two models that perform well on popular & unpopular items respectively
+LLMSeqSim &Popularity 6.2 60 60 60 Semantic item recommendation model enriched with popularity
+Table 3.
+The recommendation models used in the experiments including references to the section where they are described, the number of trials carried out for each dataset (BeauTy, Delivery Hero, STeam) per model in the hyperparameter tuning process, and a short description of each model.
 7.1.4. Hyperparameter Tuning
-We systematically tuned all models (except LLMSeqPrompt) on three validation folds with the Tree Parzen Estimator (TPE) sampler (Bergstra et al., 2011), and the total NDCG@20 across the folds as the optimization goal.
+We systematically tuned all models (except LLMSeqPrompt, LlamaSeqPromptGenItem, S3-Rec, and TALLRec, which we discuss separately) on three validation folds with the Tree Parzen Estimator (TPE) sampler (Bergstra et al., 2011), and the total NDCG@20 across the folds as the optimization goal.
 We chose to optimize for 20 recommendations to provide enough room for the models to manifest their performance differences, although top-10 and top-20 metrics are usually closely related.
@@
 Finally, the hyperparameter tuning of the hybrids involves the following main parameters: embedding source, embedding dimensionality reduction, sequential model (for the LLMSeqSim &Sequential hybrid), and popularity threshold value.
+7.1.5. Training and Evaluation of S3-Rec
+The S3-Rec model is different in nature from other models, because it is a pre-training based approach.
+Nonetheless, we were interested how this particular model would fare in a comparison with the other models in our study.
+In terms of datasets, we included experiments only for the Beauty and Steam datasets, as these include metadata about item attributes that are required to pre-train S3-Rec.
+Using the original code of S3-Rec provided by the authors, we pre-trained S3-Rec for 10 epochs, fine-tuned it for 200 epochs with early stopping enabled.
+As for the other models in our comparison, we used the NDCG and the hit rate as evaluation measures.
+We note that we determined the metric values using the evaluation protocol from the original paper.
+S3-Rec’s evaluation protocol involves using each session’s second-to-last item for validation during fine-tuning.
+The last item of each session is used for testing.
+We note that the evaluation protocol (and code) used for the other models, which is more common in the literature, is slightly different and more challenging for the algorithms, as we hide entire sessions from the validation and training data.
+As a result, the results reported for the S3-Rec model are expected to be higher than what we would observe if the exact same protocol would be applied.
+As our results will show, however, the accuracy results obtained with S3-Rec fall behind the other models, even though more information was available than for the other models.
+7.1.6. Training and Evaluation of TALLRec
+For TALLRec, we present experiments only for the Beauty dataset because the model would require many weeks of execution on the Delivery Hero and Steam datasets.
+Regarding the model’s training, we used the same preprocessing and train-validation-test split for Beauty as for LLMSeqPromptGenItem (Section 4.1).
+To make TALLRec comparable to our work, we converted TALLRec’s binary classification problem into a recommendation generation problem.
+We also used the same system prompt, user prompt template, and assistant prompt template as in LLMSeqPromptGenItem but we adapted it to TALLRec’s format.
+[^23] https://github.com/SAI990323/TALLRec/blob/main/finetune_rec.py#L301 We then trained TALLRec on an EC2 g5.12xlarge instance with 4x24GB A10G GPUs using the default hyperparameter setup provided by the authors.
+[^24] https://github.com/SAI990323/TALLRec/blob/main/shell/instruct_7B.sh
+For the evaluation, we used the same protocol as for LLMSeqPromptGenItem.
+After resolving out-of-memory issues, we asked the model to produce top-20 recommendations for each test case by repeating each test prompt 20 times.
+Then, we mapped each recommendation of the model that is not in the product catalog to the closest catalog product via OpenAI embeddings.
+Finally, we computed the evaluation metrics that we report in the paper.
+7.1.7. Fine-tuning and Evaluation of LlamaSeqPromptGenItem
+We fine-tuned LlamaSeqPromptGenItem, specifically meta-textgeneration-llama-3-8b, on the Beauty and Steam datasets using the LoRA approach for five epochs with 0.0001 learning rate, 32 alpha, 0.05 dropout, and we set the r parameter to 8.
+In addition, we used 20% of the dataset as validation set.
+Resource constraints did not allow us to conduct experiments on the Delivery Hero dataset.
+On the same grounds, we fixed the temperature of the model to 1 and the top_p parameter to 0.25, which perform best in practice according to our experience.
 7.1.8. Model Executions on Datasets
 In principle, we ran all models on all datasets except for fine-tuning models due to budget and time limitations.
-Specifically, we applied LLMSeqPromptGenList, LLMSeqPromptClass, and LLMSeqPromptRank only on the Beauty dataset.
+Specifically, we applied LLMSeqPromptGenList, LLMSeqPromptClass, LLMSeqPromptRank, and TALLRec only on the Beauty dataset, and S3-Rec and LlamaSeqPromptGenItem only on the Beauty and Steam datasets.
 In addition, we only fine-tuned OpenAI ada on LLMSeqPromptGenItem because it is no longer available for fine-tuning.
@@
 We attribute the slightly lower performance to a potential modest instability caused by the size of the dataset and to the stochasticity of the hyperparameter search process.
-Model Sec. # Trials Description
-BT DH ST
-LLMSeqSim 3 591 620 631 Semantic item recommendation model based on OpenAI or Google embeddings
-LLMSeqPromptGenItem 4.1 16 14 4 Fine-tuned model to generate the next item
-LLMSeqPromptGenList 4.2 5 - - Fine-tuned model to generate next-item recommendations
-LLMSeqPromptClass 4.3 5 - - Fine-tuned model to classify items for next-item recommendations
-LLMSeqPromptRank 4.4 5 - - Fine-tuned model to rank next-item recommendations
-BERT4Rec 5.1 661 52 57 Bidirectional neural model based on the transformer architecture
-SASRec 5.1 413 167 146 Unidirectional neural model based on the transformer architecture
-GRU4Rec 5.1 622 168 164 Recurrent neural model based on Gated Recurrent Unit
-SKNN 5.1 553 377 291 Neighborhood-based sequential model
-LLM2BERT4Rec 5.2 1074 118 95 BERT4Rec initialized with LLM embeddings
-LLM2SASRec 5.2 1280 272 226 SASRec initialized with LLM embeddings
-LLM2GRU4Rec 5.2 1001 381 256 GRU4Rec initialized with LLM embeddings
-SKNNEmb 5.2 767 218 128 SKNN utilizing LLM embeddings to compute session similarity
-LLMSeqSim &Sequential 6.1 72 72 72 Hybrid of two models that perform well on popular & unpopular items respectively
-LLMSeqSim &Popularity 6.2 60 60 60 Semantic item recommendation model enriched with popularity
-Table 3.
-The recommendation models used in the experiments including references to the section where they are described, the number of trials carried out for each dataset (BeauTy, Delivery Hero, STeam) per model in the hyperparameter tuning process, and a short description of each model.
 7.2. Results and Discussion
@@
 LLM2GRU4Rec 0.034 0.065 0.024 0.252 0.062 11.732 0.042 0.097 0.026 0.365 0.091 11.902
+BERT2GRU4Rec 0.033 0.060 0.025 0.147 0.056 11.200 0.042 0.095 0.028 0.207 0.087 11.405
 BERT4Rec 0.033 0.064 0.023 0.108 0.061 11.789 0.041 0.096 0.025 0.161 0.089 11.937
+BERT2SASRec 0.030 0.055 0.022 0.143 0.050 11.096 0.038 0.087 0.024 0.207 0.077 11.307
 LLMSeqPromptGenItem 0.032 0.054 0.025 0.548 0.054 13.140 0.037 0.075 0.027 0.728 0.074 13.385
+BERT2BERT4Rec 0.024 0.049 0.017 0.084 0.046 11.219 0.033 0.085 0.019 0.132 0.077 11.442
 GRU4Rec 0.026 0.051 0.018 0.103 0.046 11.349 0.033 0.080 0.020 0.156 0.073 11.494
@@
 LLMSeqPromptGenList 0.029 0.053 0.022 0.703 0.053 13.587 0.033 0.069 0.023 0.831 0.069 13.652
+LlamaSeqPromptGenItem 0.023 0.041 0.018 0.499 0.041 13.366 0.028 0.061 0.019 0.669 0.060 13.451
+S3-Rec 0.020 0.041 0.027 0.067
+TALLRec 0.013 0.024 0.009 0.585 0.024 13.754 0.015 0.035 0.01 0.747 0.034 13.824
 LLMSeqPromptClass 0.006 0.012 0.005 0.016 0.010 11.041 0.008 0.017 0.005 0.016 0.012 11.072
@@
 LLMSeqSim &Sequential 0.099 0.175 0.075 0.276 0.146 10.823 0.117 0.247 0.080 0.354 0.191 11.021
+BERT2BERT4Rec 0.098 0.174 0.075 0.208 0.144 10.600 0.116 0.244 0.080 0.281 0.187 10.802
+BERT2GRU4Rec 0.098 0.175 0.075 0.260 0.147 10.766 0.116 0.245 0.080 0.333 0.189 10.963
 SKNNEmb 0.098 0.166 0.077 0.364 0.147 10.938 0.114 0.228 0.082 0.434 0.184 11.076
@@
 GRU4Rec 0.086 0.153 0.066 0.232 0.126 10.869 0.103 0.217 0.070 0.304 0.164 11.058
+BERT2SASRec 0.082 0.149 0.061 0.106 0.115 10.319 0.098 0.212 0.066 0.150 0.152 10.599
 LLMSeqPromptGenItem 0.063 0.116 0.047 0.400 0.107 12.048 0.070 0.144 0.049 0.611 0.123 13.788
@@
 LLM2BERT4Rec 0.068 0.128 0.049 0.266 0.110 9.662 0.084 0.194 0.054 0.345 0.143 9.862
+BERT2BERT4Rec 0.067 0.127 0.049 0.238 0.109 9.765 0.083 0.193 0.053 0.313 0.144 9.955
 BERT4Rec 0.066 0.126 0.047 0.242 0.106 9.583 0.083 0.194 0.052 0.317 0.141 9.813
@@
 SASRec 0.065 0.123 0.047 0.168 0.104 9.647 0.081 0.187 0.052 0.232 0.138 9.869
+BERT2GRU4Rec 0.064 0.122 0.046 0.183 0.105 9.674 0.080 0.186 0.051 0.242 0.137 9.857
+BERT2SASRec 0.060 0.116 0.043 0.106 0.099 9.527 0.076 0.178 0.048 0.142 0.129 9.738
+S3-Rec 0.057 0.111 0.072 0.173
 SKNNEmb 0.051 0.099 0.037 0.243 0.072 8.789 0.064 0.151 0.041 0.353 0.093 9.053
@@
 LLMSeqSim 0.023 0.044 0.017 0.656 0.041 11.871 0.029 0.067 0.019 0.786 0.058 11.950
+LlamaSeqPromptGenItem 0.014 0.025 0.011 0.301 0.022 11.473 0.019 0.045 0.012 0.424 0.031 11.782
 Table 6. Evaluation results for the Steam dataset
+In line with our observations regarding LLM embeddings, BERT embeddings also tend to increase the performance of neural recommendation models, but not consistently for all datasets and all models.
+In most of the cases, LLM embeddings significantly outperform BERT embeddings except for one case, for the DH dataset, where BERT embeddings elevated GRU4Rec’s performance significantly stronger than LLM embeddings.
+In summary, the experiments reveal a real potential for leveraging the semantics of a dataset’s item descriptions to uplift the performance of the recommendation models.
+This potential however depends on the dataset and different models capture the semantics in different ways and to different extents.
+Most of the time, however, embedding initialization leads to significant performance gains especially when LLM embeddings are used.
+TALLRec, S3-Rec, and LlamaSeqPromptGenItem performed average in the evaluation.
+Specifically, they trail behind the neural models for the Steam dataset and also behind the LLM generation-based approaches for the Beauty dataset.
 Furthermore, SKNNEmb outperforms SKNN on Beauty by 9% at NDCG@20 and 12% at HR@20.
@@
 We make the following observations for coverage, serendipity, and novelty.
-The LLMSeqSim model consistently leads to the best catalog coverage and novelty.
-This is not too surprising, given the nature of the approach, which is solely based on embedding similarities.
+The LLMSeqSim model consistently leads to the best catalog coverage and novelty except for Novelty@20 in the Beauty dataset where TALLRec performs better.
+The performance of LLMSeqSim in terms of catalog coverage and novelty is not too surprising, given the nature of the approach, which is solely based on embedding similarities.
 Unlike other methods that use collaborative signals, i.e., past user-item interactions, the general popularity of an item in terms of the amount of observed past interactions does not play a role in LLMSeqSim, neither directly nor implicitly.
@@
 In addition, the LLMSeqPromptGenItem task has been used to fine-tune both ada and GPT 3.5 turbo OpenAI models on the Beauty and Delivery Hero datasets.
-The results are conflicting.
+The results are inconclusive.
 On the one hand, we observe a notable performance increase of 23% in terms of NDCG@20 in favor of GPT on Beauty, but, surprisingly, on Delivery Hero ada performed significantly better than GPT for the same task achieving 18% higher NDCG@20.
 Given that the comparison scope is narrow, we can only conclude that the choice of fine-tune model depends on the dataset and that the evolution of models does not necessarily lead to higher performance.
-With respect to the temperature and top_p hyperparameters we observe that in all setups a high temperature results in lower performance as it allows for more creativity at the cost of item hallucinations.
+With respect to the temperature and top_p hyperparameters, we observe that in all setups a high temperature results in lower performance as it allows for more creativity at the cost of item hallucinations.
 For LLMSeqPromptGenItem and LLMSeqPromptGenList a low but non-zero temperature leads to the best performance whereas LLMSeqPromptClass and LLMSeqPromptRank perform best with temperature set to zero.
@@
 Dataset Task Model HR@20 NDCG@20 # Uniq. Reco. Hallucin. Emb. Sim.
-[…]
+GPT base 0.029 0.013 6553 72.38% 0.919
 GPT fntd 0.077 0.038 20056 36.11% 0.960
-[…]
+PaLM base 0.029 0.015 19330 78.63% 0.944
 LLMSeqPromptGenItem PaLM fntd 0.043 0.020 21103 56.23% 0.950
-[…]
+GPT base 0.021 0.008 28372 82.20% 0.935
 GPT fntd 0.069 0.033 47966 43.35% 0.953
-[…]
+PaLM base 0.022 0.009 40019 78.85% 0.940
 Beauty LLMSeqPromptGenList PaLM fntd 0.042 0.020 32643 21.24% 0.969
-[…]
+GPT base 0.022 0.009 41951 67.39% 0.910
 GPT fntd 0.111 0.057 55301 13.79% 0.967
-[…]
+PaLM base 0.045 0.027 50416 70.85% 0.950
 DH LLMSeqPromptGenItem PaLM fntd 0.050 0.029 43698 37.87% 0.961
-[…]
+GPT base 0.017 0.010 2534 8.41% 0.858
 GPT fntd 0.059 0.032 5701 0.36% 0.889
-[…]
+PaLM base 0.019 0.013 13303 77.05% 0.870
 Steam LLMSeqPromptGenItem PaLM fntd 0.025 0.014 382446 58.87% 0.867
@@
 Comparison between base and fine-tuned models across datasets for LLMSeqPromptGenItem and LLMSeqPromptGenList in terms of HR@20, NDCG@20, number of unique recommendations, percentage of hallucinations, and embedding similarity between hallucinations and closest match.
-Finally, to explore the learning capabilities of fine-tuned models we compare their performance to the corresponding base models.
+Finally, to explore the learning capabilities of fine-tuned models, we compare their performance to the corresponding base models.
 We observe that in all setups the fine-tuned models significantly outperform their non fine-tuned counterparts.
-For LLMSeqPromptGenItem and LLMSeqPromptGenList variants the model generates recommendations that are not part of the prompt.
+For the LLMSeqPromptGenItem and LLMSeqPromptGenList variants, the model generates recommendations that are not part of the prompt.
 For those, the base model has a much higher rate of hallucinations compared to fine-tuned models as Table 7 shows.
@@
 Therefore, the base model has to provide recommendations solely based on the information in the prompt.
-For variants LLMSeqPromptClass and LLMSeqPromptRank the recommendation options are explicitly part of the context such that the model only has to select and/or re-rank the given item names.
+For the variants LLMSeqPromptClass and LLMSeqPromptRank, the recommendation options are explicitly part of the context such that the model only has to select and/or re-rank the given item names.
 Thus, hallucinations are uncommon in those variants.
@@
 This is an expected finding that is widely acknowledged by literature and practice ( , 2023; Touvron et al., 2023).
-All in all, the fine-tuned models perform better for each recommendation task with less hallucinations and closer semantic proximity to the actual options compared to the base models.
-This finding clearly shows that fine-tuning can enable a model to learn not only a task, but also concepts of a domain to some extent and contradicts existing findings and common belief that fine-tuning is only applicable for learning form rather than facts (Mitra et al., 2023; Kirstain et al., 2022; Ovadia et al., 2024; Lewis et al., 2021; Wang et al., 2023; Kandpal et al., 2023; Luo et al., 2023; Chen et al., 2020; Kirkpatrick et al., 2017; Berglund et al., 2023).
+All in all, the fine-tuned models perform better for each recommendation task with fewer hallucinations and closer semantic proximity to the actual options compared to the base models.
+This finding clearly shows that fine-tuning can enable a model to learn not only how to perform a task, but also concepts of a domain to some extent.
+Thus, it contradicts existing findings and common belief that fine-tuning is only applicable for learning form (instructions) rather than facts (Mitra et al., 2023; Kirstain et al., 2022; Ovadia et al., 2024; Lewis et al., 2021; Wang et al., 2023; Kandpal et al., 2023; Luo et al., 2023; Chen et al., 2020; Kirkpatrick et al., 2017; Berglund et al., 2023).
 7.3.4. Hybrids
@@
 In terms of research limitations, our work so far is based on three datasets from two domains (e-commerce, games).
-As the datasets used in our research are markedly different in terms of various characteristics (e.g., sparsity, size, available meta-data), we are confident that our findings generalize beyond the conducted experiments.
+The datasets used in our research are however markedly different in terms of various characteristics (e.g., sparsity, size, available metadata, recurrence patterns, frequency of use) and in terms of their nature (e.g., one long sequence of product reviews per user for Amazon Beauty vs.
+a collection of supermarket purchases per user, each consisting of a sequence of products, in the Delivery Hero dataset).
+Therefore, we are confident that our findings generalize beyond the conducted experiments.
 Nonetheless, further experiments are needed to validate the benefits of leveraging LLMs for sequential recommendation in other domains, e.g., in music or news recommendation.
+In addition, in our current work we have not specifically analyzed the potential of LLM-enhanced sequential recommendations for cold-start situations, where very limited information is available about the preferences of individual users.
+Also, as another area for future works, it would be interesting to study the use of LLMs for cross-domain recommendation settings.
+Finally, future work might include the consideration of alternative LLMs or larger versions of existing LLMs, as these models continue to become more powerful and larger in terms of their parameters.
+One important research question in that context is if larger model sizes will have a significant impact on recommendation accuracy for the different approaches studied in this paper, see also the analyses in (Zhang et al., 2024).
 To conclude, Large Language Models and AI assistants like ChatGPT have exhibited disruptive effects in various domains in the past few years.